### v4

# Pathwise uniform convergence of a full discretization for a three-dimensional stochastic Allen-Cahn equation with multiplicative noise

Abstract
This paper analyzes a full discretization of a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. The discretization uses the Euler scheme for temporal discretization and the finite element method for spatial discretization. A key contribution of this work is the introduction of a novel stability estimate for a discrete stochastic convolution, which plays a crucial role in establishing pathwise uniform convergence estimates for fully discrete approximations of nonlinear stochastic parabolic equations. By using this stability estimate in conjunction with the discrete stochastic maximal $L^{p}$-regularity estimate, the study derives a pathwise uniform convergence rate that encompasses general spatial $L^{q}$-norms. Moreover, the theoretical convergence rate is verified by numerical experiments.
Keywords: stochastic Allen-Cahn equation, Euler scheme, the finite element method, discrete stochastic maximal $L^{p}$-regularity
1 Introduction
Assume that $(\Omega,\mathcal{F},\mathbb{P})$ is a given complete probability space with a filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ are defined. Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$. For each $t\in\mathbb{R}_{+}$, define $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ by
$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H.$
Let $\mathcal{O}$ be a bounded convex polyhedral domain in $\mathbb{R}^{3}$, and let $\Delta$ be the Laplace operator with homogeneous Dirichlet boundary condition. This paper considers the following stochastic Allen-Cahn equation:
$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0<t\leqslant T,\\ y(0)=v,\end{cases}$ (1.1)
where $T\in(0,\infty)$ and the initial value $v$ are given, and the operator $F$ will be defined later.
The Allen-Cahn equation is an important mathematical model in materials science [1] and it is also of fundamental importance for the geometric moving interface problem [6]. By incorporating thermal fluctuations, impurities, and atomistic processes that describe surface motion, the stochastic Allen-Cahn equation naturally arises [8, 17, 31].
Recently, there has been a growing interest in the numerical analysis of the stochastic Allen-Cahn equation. A brief summary of some recent works is given below. For the stochastic Allen-Cahn equation with additive noise, Becker and Jentzen [3] obtained strong convergence rates for a nonlinearity-truncated Euler-type approximation. Becker et al. [2] established sharp strong convergence rates for an explicit space-time discrete numerical approximation. Bréhier et al. [4] analyzed an explicit temporal splitting numerical scheme. Qi and Wang [29] derived strong error estimates for a full discretization using the finite element method for spatial discretization and the Euler scheme for temporal discretization. In addition, Wang [37] analyzed an efficient explicit full-discrete scheme for a one-dimensional stochastic Allen–Cahn equation. The numerical analysis of the stochastic Allen-Cahn equation with multiplicative noise generally presents greater challenges. In this regard, Feng et al. [7] derived the strong convergence of finite element approximations of a stochastic Allen-Cahn equation with finite-dimensional gradient-type multiplicative noise, Majee and Prohl [27] analyzed a space-time discretization, and Liu and Qiao [26] investigated the strong convergence of Galerkin-based Euler approximations. We also refer the reader to [5, 10, 11, 13, 14, 15, 18, 19, 30, 32] and the references therein for more related works.
So far, the numerical analysis of the stochastic Allen-Cahn equation in the literature generally considers only the $L^{p}(\Omega;L^{2})$-norm error estimate for the fully discrete approximations. There is a lack of pathwise uniform convergence estimates and convergence estimates with the general spatial $L^{q}$-norms. It is worth noting that Bréhier et al. [4] derived some pathwise uniform convergence estimates for an explicit temporal splitting numerical scheme; however, their analysis may not be directly applicable to general convex bounded domains (see [4, pp. 2116]), and extending this analysis to the full discretization of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise is nontrivial.
The main contributions of this paper are summarized as follows:
  • We establish the stability estimate
    $\mathbb{E}\max_{1\leqslant j\leqslant J}\Big{\lVert}{\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}(I-\tau\Delta_{h})^{k-j}P_{h}g(t)\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{q}}^{p}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}$
    for all $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$ with $p\in(2,\infty)$ and $q\in[2,\infty)$. This estimate is significant as it facilitates the demonstration of the pathwise uniform convergence of fully discrete approximations for nonlinear stochastic parabolic equations. While Van Neerven and Veraar [36] have established similar estimates for the Euler scheme, their approach is not directly applicable to deriving the above stability estimate, as detailed in Section 3.2.
  • For a fully discrete approximation of Eq. 1.1, using the Euler scheme for temporal discretization and the finite element method for spatial discretization, we establish the following error estimate:
    $\Big{\lVert}{\max_{1\leqslant j\leqslant J}\lVert{y(t_{j})-Y_{j}}\rVert_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}\leqslant c\left(h^{2-\varepsilon}+\tau^{1/2}\right)$
    for all $p\in(2,\infty)$, $q\in[2,\infty)$ and $\varepsilon>0$. Notably, even for the three-dimensional stochastic Allen-Cahn equation with additive noise, to our knowledge, analogous error estimates for fully discrete approximations utilizing the Euler scheme have yet to be discovered. This finding fills an important gap in the existing literature.
Our analysis builds upon the discrete stochastic maximal $L^{p}$-regularity estimate developed in [21, 23], which was inspired by the stochastic maximal $L^{p}$-regularity theory in [35]. A key role is also played by a stability estimate of a spatial semi-discretization from our prior work [22]. We impose relatively mild regularity conditions on the operator $F$ in Eq. 1.1, and although our analysis assumes that the initial value $v$ belongs to $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, it can be extended to cases where the initial value $v$ has lower regularity. Our work introduces a novel method for the numerical analysis of nonlinear stochastic parabolic equations, using the discrete stochastic maximal $L^{p}$-regularity estimate. This offers a fresh perspective compared to previous studies in the field.
The remainder of this paper is organized as follows. In Section 2, we describe a full discretization of Eq. 1.1 and present a pathwise uniform convergence estimate. In Section 3, we establish a stability estimate for a discrete stochastic convolution, and prove the convergence estimate discussed earlier. Section 4 presents numerical experiments to validate the established pathwise uniform convergence. Finally, Section 5 provides a summary of the findings and contributions of this study.
2 Full discretization
We first introduce some notations. For any Banach space $X$, let $\gamma(H,X)$ denote the space of all $\gamma$-radonifying operators from $H$ to $X$; see [12, Chapter 9]. For any $p\in(1,\infty)$ and Banach space $X$, let $L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;X))$ and $L_{\mathbb{F}}^{p}(\Omega\times(0,T);X)$ denote the spaces of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega;L^{2}(0,T;X))$ and $L^{p}(\Omega\times(0,T);X)$, respectively. For ease of notation, we abbreviate the Lebesgue space $L^{q}(\mathcal{O})$ to $L^{q}$ for all $q\in[1,\infty]$. In addition, let $W_{0}^{1,\infty}(\mathcal{O})$ and $W^{2,\infty}(\mathcal{O})$ denote two standard Sobolev spaces; see, e.g., [33].
Next, we define the operator $F$ in Eq. 1.1 as follows. Let $(f_{n})_{n\in\mathbb{N}}$ denote a sequence of continuous functions from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$ such that for each $n\in\mathbb{N}$,
$\begin{cases}f_{n}(x,0)=0,&\forall x\in\partial\mathcal{O},\\ \lvert{f_{n}(x,y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert),&\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},\\ \lvert{f_{n}(x,y_{1})-f_{n}(x,y_{2})}\rvert\leqslant C_{F}\lvert{y_{1}-y_{2}}\rvert,&\forall x\in\mathcal{O},\,\forall y_{1},y_{2}\in\mathbb{R},\\ \lvert{f_{n}(x_{1},y)-f_{n}(x_{2},y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert)\lvert{x_{1}-x_{2}}\rvert,&\forall x_{1},x_{2}\in\mathcal{O},\,\forall y\in\mathbb{R},\end{cases}$ (2.1)
where $C_{F}$ is a positive constant independent of $n$, $x$, $y$, $x_{1}$, $x_{2}$, $y_{1}$ and $y_{2}$. With a little abuse of notation, for each $n\in\mathbb{N}$ we will use $f_{n}$ to denote a Nemytski operator defined by $f_{n}(v):=f_{n}(\cdot,v(\cdot))$, $\forall v\in L^{2}$. Let $\{\lambda_{n}\}_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying $\sum_{n\in\mathbb{N}}\lambda_{n}<\infty$. Finally, for any $v\in L^{2}$ define
$F(v):=\sum_{n=0}^{\infty}\sqrt{\lambda}_{n}h_{n}\otimes f_{n}(v),$
where $h_{n}\otimes f_{n}(v)\in\gamma(H,L^{2})$ is defined by
$\big{(}h_{n}\otimes f_{n}(v)\big{)}(w):=(w,h_{n})_{H}f_{n}(v),\quad\forall w\in H.$
By [12, Proposition 9.1.9] and (2.1), a straightforward calculation gives
$\lVert{F(v)}\rVert_{\gamma(H,L^{2})}\leqslant c\big{(}1+\lVert{v}\rVert_{L^{2}}\big{)},\quad\forall v\in L^{2},$
where $c$ is a positive constant depending only on $C_{F}$ in (2.1), $\sum_{n=0}^{\infty}\lambda_{n}$, and the volume of $\mathcal{O}$.
Now, we introduce the full discretization for Eq. 1.1. Let $J$ be a positive integer, and set $t_{j}:=j\tau$ for each $0\leqslant j\leqslant J$, where $\tau:=T/J$. Let $\mathcal{K}_{h}$ be a conforming and quasi-uniform triangulation of $\mathcal{O}$ consisting of three-dimensional simplexes, and let $h$ denote the maximum size of the elements in $\mathcal{K}_{h}$. Define
$X_{h}:=\left\{v_{h}\in C(\overline{\mathcal{O}})\;\middle|\;\text{ $v_{h}$ is linear on each $K\in\mathcal{K}_{h}$ and $v_{h}$ vanishes on $\partial\mathcal{O}$}\right\},$
where $C(\overline{\mathcal{O}})$ is the set of all continuous functions on $\overline{\mathcal{O}}$. Let $P_{h}$ be the $L^{2}$-orthogonal projection onto $X_{h}$, and for any $u_{h}\in X_{h}$ define $\Delta_{h}u_{h}\in X_{h}$ by
$\int_{\mathcal{O}}(\Delta_{h}u_{h})\cdot v_{h}\,\mathrm{d}\mu=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}\mu,\quad\forall v_{h}\in X_{h},$
where $\mu$ denotes the three-dimensional Lebesgue measure on $\mathcal{O}$. The full discretization of Eq. 1.1 is given as follows:
[…]
Finally, we establish the convergence of the discretization Eq. 2.2 in the following theorem; the proof is deferred to Section 3.
Theorem 2.1.
Assume that $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ and $\tau\leqslant h^{2}$. Let $y$ and $(Y_{j})_{j=0}^{J}$ be the solutions of Eqs. 1.1 and 2.2, respectively. Then for all $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following error estimate holds:
$\Big{\lVert}{\max_{1\leqslant j\leqslant J}\lVert{Y_{j}-y(t_{j})}\rVert_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}\leqslant c\big{(}h^{2-\varepsilon}+\tau^{1/2}\big{)},$ (2.3)
where $c$ denotes a positive constant independent of the spatial mesh size $h$ and the time step $\tau$.
Remark 2.1.
Regrading the regularity of the solution $y$ of Eq. 1.1, we refer the reader to [24, 25, 26], where the authors have derived some regularity estimates with the spatial $L^{2}$-norm and the spatial $H^{1}$-norm. Furthermore, the authors in [22] have extended these results by establishing regularity estimates in terms of the general spatial $L^{q}$-norms; for example, under the condition $v\in L^{\infty}$ they have shown that
$y\in L^{p}(\Omega;C([0,T];L^{q})),\quad\forall p,q\in(2,\infty).$
Remark 2.2.
For Theorem 2.1, we have the following comments.
  1. 1.
    The condition $\tau\leqslant h^{2}$ can be relaxed by re-examining and refining the analysis in Subsection 3.3.3.
  2. 2.
    The spatial accuracy of $O(h^{2-\epsilon})$ in Eq. 2.3 can be improved to $O(h^{2})$ by imposing stronger regularity conditions on the operator $F$. Specifically, if there exist an $\epsilon\in(0,1)$ and a positive constant $c$ such that
    $\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1+\epsilon,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{\dot{H}^{1+\epsilon,q}}\big{)}$
    for all $v\in\dot{H}^{1+\epsilon,q}$ with $q\in[2,\infty)$, then the improved accuracy can be achieved. The above space $\dot{H}^{1+\epsilon,q}$ is defined in the subsequent section.
  3. 3.
    The authors in [22] derived a strong convergence rate for a spatial semi-discretization of Eq. 1.1 with a rough initial value $v\in L^{\infty}$. However, to the best of our knowledge, there is currently no strong convergence estimate for fully discrete approximations of the three-dimensional stochastic Allen-Cahn equation with a rough initial value. It would be of interest to establish similar error estimates as Eq. 2.3 for fully discrete approximations with rough initial values.
3 Proofs
In this section, we adopt the following notational conventions:
  • The symbol $\mathbb{E}$ represents the expectation operator associated with the probability space $(\Omega,\mathcal{F},\mathbb{P})$.
  • For any two Banach spaces $E_{1}$ and $E_{2}$, $\mathcal{L}(E_{1},E_{2})$ denotes the space of all bounded linear operators from $E_{1}$ to $E_{2}$, and $\mathcal{L}(E_{1},E_{1})$ is abbreviated to $\mathcal{L}(E_{1})$.
  • The operator $I$ denotes the identity map.
[…]
  • For $\alpha\in\mathbb{R}$ and $q\in(1,\infty)$, $\dot{H}_{h}^{\alpha,q}$ is defined as the Banach space $X_{h}$ equipped with the norm
    $\lVert{v_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}}:=\lVert{(-\Delta_{h})^{\alpha/2}v_{h}}\rVert_{L^{q}},\quad\forall v_{h}\in X_{h}.$
  • For $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, equipped with the norm
    $\lVert{v}\rVert_{\dot{H}^{\alpha,q}}:=\lVert{(-\Delta)^{\alpha/2}v}\rVert_{L^{q}},\quad\forall v\in\dot{H}^{\alpha,q}.$
  • Specifically, $\dot{H}^{-1,q}$ denotes the dual space of $\dot{H}^{1,q/(q-1)}$ for each $q\in(1,\infty)$.
  • The symbol $c$ is used to denote a generic positive constant, which is independent of $h$ and $\tau$, and its specific value may change from line to line.
The subsequent structure of this section is outlined as follows. In Subsection 3.1, we introduce several preliminary lemmas that form the foundation for our subsequent analysis. Subsequently, Subsection 3.2 is dedicated to establishing a stability estimate for a discrete stochastic convolution, which is of independent interest and also vital for the proof of Theorem 2.1. Lastly, Subsection 3.3 presents a detailed proof of Theorem 2.1.
3.1 Preliminary lemmas
Lemma 3.1.
For any $\alpha\in[0,2]$ and $q\in(1,\infty)$, $\lVert{P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},\,\dot{H}_{h}^{\alpha,q})}$ is uniformly bounded with respect to (w.r.t.) the spatial mesh size $h$.
Lemma 3.2.
Let $q\in(1,\infty)$ and $0\leqslant\alpha\leqslant\beta<\infty$. Then the following holds:
$\lVert{e^{t\Delta_{h}}}\rVert_{\mathcal{L}(\dot{H}_{h}^{\alpha,q},\dot{H}_{h}^{\beta,q})}\leqslant ct^{(\alpha-\beta)/2},\quad\forall t\in(0,T];$ (3.1)
$\lVert{I-e^{t\Delta_{h}}}\rVert_{\mathcal{L}(\dot{H}_{h}^{\beta,q},\dot{H}_{h}^{\alpha,q})}\leqslant ct^{(\beta-\alpha)/2},\quad\forall t\in[0,T];$ (3.2)
$\lVert{(I-\tau\Delta_{h})^{-1}-I}\rVert_{\mathcal{L}(\dot{H}_{h}^{\beta,q},\dot{H}_{h}^{\alpha,q})}\leqslant c\tau^{(\beta-\alpha)/2};$ (3.3)
$\lVert{(I-\tau\Delta_{h})^{-m}}\rVert_{\mathcal{L}(\dot{H}_{h}^{\alpha,q},\dot{H}_{h}^{\beta,q})}\leqslant\frac{c}{(m\tau)^{(\beta-\alpha)/2}},\quad\,\forall m\in\mathbb{N}_{>0}.$ (3.4)
Lemma 3.3.
Let $p\in(1,\infty)$ and $q\in[2,\infty)$. There exists a positive constant $c$ such that for any $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$, the following inequality holds:
$\Big{\lVert}{\int_{0}^{T}g(t)\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{p}(\Omega;L^{q})}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega;L^{2}(0,T;\gamma(H;L^{q})))}.$
Lemma 3.4.
Let $q\in[2,\infty)$. Then the following holds:
$\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),$ $\quad\forall v\in L^{q},$ (3.5)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}),$ $\forall v\in\dot{H}^{1,q},$ (3.6)
$\lVert{F(u)-F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},$ $\forall u,v\in L^{q}.$ (3.7)
Lemma 3.5.
Let $p,q\in(1,\infty)$. Assume that $(Z_{j})_{j=0}^{J}$ is the solution of the discretization
$Z_{j+1}-Z_{j}=\tau\Delta_{h}Z_{j+1}+P_{h}\int_{t_{j}}^{t_{j+1}}g(t)\,\mathrm{d}t,\quad 0\leqslant j<J,$
$Z_{0}=0,$
where $g\in L^{p}(0,T;L^{q})$. Then
$\Big{(}\tau\sum_{j=1}^{J}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{2,q}}^{p}\Big{)}^{1/p}\leqslant c\lVert{g}\rVert_{L^{p}(0,T;L^{q})}.$
Remark 3.1.
Lemmas 3.1 and 3.2 are well known; see, e.g., [34]. For the proof of Lemma 3.3, we refer the reader to, e.g., [35, pp. 792]. Lemma 3.4 is contained in [22, Lemma 4.9]. For the proof of Lemma 3.5, we refer the reader to [16, 20].
3.2 A stability estimate for a discrete stochastic convolution
The objective of this subsection is to establish the stability estimate
$\bigg{(}\mathbb{E}\Big{[}\max_{1\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}\bigg{)}^{1/p}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$ (3.8)
for the discrete stochastic convolution defined by
$Z_{j}=\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}(I-\tau\Delta_{h})^{k-j}P_{h}g(t)\,\mathrm{d}W_{H}(t),\quad 1\leqslant j\leqslant J,$ (3.9)
where $g$ belongs to $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, with parameters $p\in(2,\infty)$ and $q\in[2,\infty)$. This stability estimate is crucial for the derivation of pathwise uniform convergence rates for the numerical approximations of a broad class of nonlinear stochastic parabolic equations.
Within the framework of Hilbert spaces, we note that Gyöngy and Millet [9, Theorem 4.4] have established the following stability estimate:
$\left(\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\|Z_{j}\|_{L^{2}}^{2}\right]\right)^{1/2}\leqslant c\left\|g\right\|_{L^{2}(\Omega\times(0,T);\gamma(H,L^{2}))}.$
Moreover, under the Banach space setting, a pivotal advancement was made by van Neerven and Veraar [36, Proposition 5.4], who demonstrated the following sharp estimate for a discrete stochastic convolution without spatial discretization:
$\left(\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\Big{\|}\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}(I-\tau\Delta)^{k-j}g(t)\,\mathrm{d}W_{H}(t)\Big{\|}_{L^{q}}^{p}\right]\right)^{1/p}\leqslant c\left\|g\right\|_{L^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))},$
valid for all $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$ with $p,q\in[2,\infty)$. The derivation of this estimate relies on the contraction property of the operator $(I-\tau\Delta)^{-1}$ on $L^{q}$ for all $q\in[2,\infty)$. However, it is important to note that $(I-\tau\Delta_{h})^{-1}$ fails to possess the contraction property on $\dot{H}_{h}^{0,q}$ for any $q\in(2,\infty)$. Hence, Proposition 5.4 from [35] can not be directly applied to bound the discrete stochastic convolution defined by Eq. 3.9.
To address the aforementioned challenge, we employ the approximation property of $(I-\tau\Delta_{h})^{-1}$ to the identity operator $I$ as presented in Eq. 3.3, in conjunction with the discrete stochastic maximal $L^{p}$-regularity estimate for
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p},\quad p\in(2,\infty),\,q\in[2,\infty).$
We formally state the discrete stochastic maximal $L^{p}$-regularity estimate in the following proposition.
Proposition 3.1.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Assume that $(Z_{j})_{j=1}^{J}$ is defined by Eq. 3.9, where $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$. Then the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
$\Big{(}\tau\sum_{j=1}^{J}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\Big{)}^{1/p}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}.$
Proof.
Note that [23, Theorem 3.1] has demonstrated that the $H^{\infty}$-calculus of $-\Delta_{h}$ is uniformly bounded with respect to the spatial mesh size $h$. Subsequently, by closely following the arguments outlined in the proof of [21, Theorem 3.2], we establish the desired discrete stochastic maximal $L^{p}$-regularity estimate. ∎
Now we present the main result of this subsection in the form of the following proposition.
Proposition 3.2.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Suppose that $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, and $(Z_{j})_{j=1}^{J}$ are defined by Eq. 3.9. Then the stability estimate Eq. 3.8 holds.
Proof.
Let $Z_{0}:=0$ and, for any $t\in[0,T]$, define
$\Psi(t):=(I-\tau\Delta_{h})^{-1}P_{h}g(t),\quad\mathcal{G}(t):=p\sum_{j=0}^{J-1}\mathbbm{1}_{[t_{j},t_{j+1})}(t)\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\left\langle\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\Psi(t)\right\rangle,$
where $\mathbbm{1}_{[t_{j},t_{j+1})}$ is the indicator function for the time interval $[t_{j},t_{j+1})$. For any $0\leqslant j<J$, we also define
$\delta_{j}:=(I-\tau\Delta_{h})^{-1}Z_{j}-Z_{j},\quad M_{j}:=\delta_{j}+\int_{t_{j}}^{t_{j+1}}\Psi(t)\,\mathrm{d}W_{H}(t).$
It can be verified that
$Z_{j+1}=Z_{j}+M_{j}\quad\text{for all $0\leqslant j<J$}.$
The rest of the proof is divided into the following five steps.
Step 1. Let us prove
$\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}\leqslant I_{1}+I_{2}+I_{3}+I_{4},$ (3.10)
where
$\hskip-165.00024ptI_{1}:=p\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\left|\left\langle\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\delta_{j}\right\rangle\right|\bigg{]},$
$\hskip-195.00029ptI_{2}:=\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\Big{\lvert}{\int_{0}^{t}\mathcal{G}(t)\,\mathrm{d}W_{H}(t)}\Big{\rvert}\bigg{]},$
$I_{3}:=p(p-q)\sum_{j=0}^{J-1}\mathbb{E}\bigg{[}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2q}\Big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\Big{\rangle}^{2}\,\mathrm{d}\theta\bigg{]},$
$\hskip-60.00009ptI_{4}:=p(q-1)\sum_{j=0}^{J-1}\mathbb{E}\bigg{[}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2},\,M_{j}^{2}}\big{\rangle}\,\mathrm{d}\theta\bigg{]}.$
Fix any $0\leqslant j<J$. Define the function $g:[0,1]\to\mathbb{R}$ by
$g(\theta):=\big{\lVert}{Z_{j}+\theta M_{j}}\big{\rVert}_{L^{q}}^{p},\quad\theta\in[0,1].$
A straightforward computation yields
$g^{\prime}(0)=p\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,M_{j}}\big{\rangle}$
$\qquad=p\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\delta_{j}}\big{\rangle}+\int_{t_{j}}^{t_{j+1}}\mathcal{G}(t)\,\mathrm{d}W_{H}(t)$
and
$g^{\prime\prime}(\theta)=p(p-q)\big{\lVert}{Z_{j}+\theta M_{j}}\big{\rVert}_{L^{q}}^{p-2q}\Big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\Big{\rangle}^{2}+{}$
$\qquad\qquad\quad p(q-1)\big{\lVert}{Z_{j}+\theta M_{j}}\big{\rVert}_{L^{q}}^{p-q}\Big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2},\,M_{j}^{2}}\Big{\rangle},\quad\forall\theta\in[0,1].$
Hence, by $g(1)=g(0)+g^{\prime}(0)+\int_{0}^{1}(1-\theta)g^{\prime\prime}(\theta)\mathrm{d}\theta$ and the fact that $g(1)=\lVert{Z_{j+1}}\rVert_{L^{q}}^{p}$ and $g(0)=\lVert{Z_{j}}\rVert_{L^{q}}^{p}$, we obtain
$\lVert{Z_{j+1}}\rVert_{L^{q}}^{p}$ $\leqslant\lVert{Z_{j}}\rVert_{L^{q}}^{p}+p\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\delta_{j}}\big{\rangle}+\int_{t_{j}}^{t_{j+1}}\mathcal{G}(t)\,\mathrm{d}W_{H}(t)$
$\quad{}+p(p-q)\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\big{\rangle}^{2}\,\mathrm{d}\theta$
$\quad{}+p(q-1)\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2},\,M_{j}^{2}}\big{\rangle}\,\mathrm{d}\theta.$
Given that $Z_{0}=0$ and the preceding inequality holds for all $0\leqslant j<J$, we deduce the following bound for the supremum of the sequence $(Z_{j})_{j=0}^{J}$:
$\max_{0\leqslant j\leqslant J}\left\|Z_{j}\right\|_{L^{q}}^{p}\leqslant p\sum_{j=0}^{J-1}\left\|Z_{j}\right\|_{L^{q}}^{p-q}\left|\left\langle\left|Z_{j}\right|^{q-2}Z_{j},\delta_{j}\right\rangle\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\mathcal{G}(s)\,\mathrm{d}W_{H}(s)\right|$
$\qquad\qquad{}+p(p-q)\sum_{j=0}^{J-1}\int_{0}^{1}\left\|Z_{j}+\theta M_{j}\right\|_{L^{q}}^{p-2q}\left\langle\left|Z_{j}+\theta M_{j}\right|^{q-2}(Z_{j}+\theta M_{j}),M_{j}\right\rangle^{2}\,\mathrm{d}\theta$
$\qquad\qquad{}+p(q-1)\sum_{j=0}^{J-1}\int_{0}^{1}\left\|Z_{j}+\theta M_{j}\right\|_{L^{q}}^{p-q}\left\langle\left|Z_{j}+\theta M_{j}\right|^{q-2},M_{j}^{2}\right\rangle\,\mathrm{d}\theta,$
which directly implies the desired inequality (3.10) upon taking expectations on both sides.
Step 2. Let us prove
$I_{1}\leqslant c\tau\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\lVert{Z_{j}}\rVert_{L^{q}}^{p}+\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\Big{]}.$ (3.11)
From the definition of $I_{1}$, it follows that
$I_{1}$ $\leqslant p\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\lVert}{\lvert{Z_{j}}\rvert^{q-2}Z_{j}}\big{\rVert}_{\dot{H}^{1,q/(q-1)}}\lVert{\delta_{j}}\rVert_{\dot{H}^{-1,q}}\Big{]}$
$\leqslant c\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\lVert}{\lvert{Z_{j}}\rvert^{q-2}Z_{j}}\big{\rVert}_{\dot{H}^{1,q/(q-1)}}\lVert{\delta_{j}}\rVert_{\dot{H}_{h}^{-1,q}}\Big{]}.$
For any $0\leqslant j<J$, we observe that
$\|\delta_{j}\|_{\dot{H}_{h}^{-1,q}}$ $\leqslant\|(I-\tau\Delta_{h})^{-1}-I\|_{\mathcal{L}(\dot{H}_{h}^{1,q},\dot{H}_{h}^{-1,q})}\|Z_{j}\|_{\dot{H}_{h}^{1,q}}$
$=\|(I-\tau\Delta_{h})^{-1}-I\|_{\mathcal{L}(\dot{H}_{h}^{2,q},\dot{H}_{h}^{0,q})}\|Z_{j}\|_{\dot{H}_{h}^{1,q}}$
$\leqslant c\tau\|Z_{j}\|_{\dot{H}_{h}^{1,q}},$
the last inequality justified by (3.3) with $\alpha=0$ and $\beta=2$. Furthermore,
$\|\,|Z_{j}|^{q-2}Z_{j}\|_{\dot{H}^{1,q/(q-1)}}$ $\leqslant c\|\nabla(|Z_{j}|^{q-2}Z_{j})\|_{L^{q/(q-1)}}$
[…]
$\leqslant c\|Z_{j}\|_{L^{q}}^{q-2}\|Z_{j}\|_{\dot{H}_{h}^{1,q}}.$
Consequently, we arrive at
$I_{1}\leqslant c\tau\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\|Z_{j}\|_{L^{q}}^{p-2}\|Z_{j}\|_{\dot{H}_{h}^{1,q}}^{2}\Big{]},$
which leads to Eq. 3.11 by Young’s inequality.
Step 3. Let us prove
$I_{2}\leqslant\varepsilon\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}+c\varepsilon^{1-p}\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},\quad\forall\varepsilon\in(0,1).$ (3.12)
By the Burkholder-Davis-Gundy inequality and Hölder’s inequality, we obtain
$I_{2}$ $\leqslant c\mathbb{E}\Big{[}\int_{0}^{T}\lVert{\mathcal{G}(t)}\rVert_{\gamma(H,\mathbb{R})}^{2}\,\mathrm{d}t\Big{]}^{1/2}$
$=c\mathbb{E}\Big{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\big{\lVert}{\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\Psi(t)}\big{\rangle}}\big{\rVert}_{\gamma(H,\mathbb{R})}^{2}\,\mathrm{d}t\Big{]}^{1/2}$
$\leqslant c\mathbb{E}\Big{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{Z_{j}}\rVert_{L^{q}}^{2p-2q}\big{\lVert}{\lvert{Z_{j}}\rvert^{q-2}Z_{j}}\big{\rVert}_{L^{q/(q-1)}}^{2}\lVert{\Psi(t)}\rVert_{\gamma(H,L^{q})}^{2}\,\mathrm{d}t\Big{]}^{1/2}$
$=c\mathbb{E}\Big{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{Z_{j}}\rVert_{L^{q}}^{2p-2}\lVert{\Psi(t)}\rVert_{\gamma(H,L^{q})}^{2}\,\mathrm{d}t\Big{]}^{1/2}.$
For any $\varepsilon\in(0,1)$, we can further bound $I_{2}$ as follows:
$I_{2}$ $\leqslant c\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p-1}\lVert{\Psi}\rVert_{L^{2}(0,T;\gamma(H,L^{q}))}\Big{]}$
$\leqslant c\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p-1}\lVert{\Psi}\rVert_{L^{p}(0,T;\gamma(H,L^{q}))}\Big{]}$
$\leqslant\varepsilon\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}+c\varepsilon^{1-p}\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},$
where the second inequality uses Hölder’s inequality, and the third inequality applies Young’s inequality. This proves the desired inequality Eq. 3.12.
Step 4. Let us prove that
$I_{3}+I_{4}\leqslant c\tau\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\lVert{Z_{j}}\rVert_{L^{q}}^{p}+\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\Big{]}+c\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$ (3.13)
To this end, we apply Hölder’s inequality and Young’s inequality accordingly:
$I_{3}+I_{4}$ $\leqslant c\sum_{j=0}^{J-1}\mathbb{E}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2}\lVert{M_{j}}\rVert_{L^{q}}^{2}\,\mathrm{d}\theta$
$\leqslant c\sum_{j=0}^{J-1}\mathbb{E}\int_{0}^{1}\tau\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p}+\tau^{1-p/2}\lVert{M_{j}}\rVert_{L^{q}}^{p}\,\mathrm{d}\theta$
$\leqslant c\tau\sum_{j=0}^{J-1}\mathbb{E}\lVert{Z_{j}}\rVert_{L^{q}}^{p}+c\tau^{1-p/2}\sum_{j=0}^{J-1}\mathbb{E}\lVert{M_{j}}\rVert_{L^{q}}^{p}.$
On the other hand,
$\sum_{j=0}^{J-1}\mathbb{E}\lVert{M_{j}}\rVert_{L^{q}}^{p}$ $\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\sum_{j=0}^{J-1}\mathbb{E}\lVert{\delta_{j}}\rVert_{L^{q}}^{p}+c\sum_{j=0}^{J-1}\mathbb{E}\big{\lVert}{\int_{t_{j}}^{t_{j+1}}\Psi(t)\,\mathrm{d}W_{H}(t)}\big{\rVert}_{L^{q}}^{p}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\sum_{j=0}^{J-1}\mathbb{E}\lVert{\delta_{j}}\rVert_{L^{q}}^{p}+c\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\int_{t_{j}}^{t_{j+1}}\lVert{\Psi(t)}\rVert_{\gamma(H,L^{q})}^{2}\,\mathrm{d}t\Big{]}^{p/2}$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}c\sum_{j=0}^{J-1}\mathbb{E}\lVert{\delta_{j}}\rVert_{L^{q}}^{p}+c\sum_{j=0}^{J-1}\tau^{p/2-1}\int_{t_{j}}^{t_{j+1}}\lVert{\Psi(t)}\rVert_{L^{p}(\Omega;\gamma(H,L^{q}))}^{p}\,\mathrm{d}t$
$\stackrel{{\scriptstyle\mathrm{(iv)}}}{{\leqslant}}c\tau^{p/2}\sum_{j=0}^{J-1}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}+c\tau^{p/2-1}\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},$
where in (i) we use the definition $M_{j}:=\delta_{j}+\int_{t_{j}}^{t_{j+1}}\Psi(t)\,\mathrm{d}W_{H}(t)$, in (ii) we invoke Lemma 3.3, in (iii) we apply Hölder’s inequality, and in (iv) we utilize the definition $\delta_{j}:=(I-\tau\Delta_{h})^{-1}Z_{j}-Z_{j}$, along with Eq. 3.3 with $\alpha=0$ and $\beta=1$. The desired inequality Eq. 3.13 then follows by combining the above estimates.
Step 5. Combining Eqs. 3.10, 3.11, 3.12 and 3.13 yields
$\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}$ $\leqslant\varepsilon\mathbb{E}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}+c\tau\sum_{j=0}^{J-1}\mathbb{E}\Big{(}\lVert{Z_{j}}\rVert_{L^{q}}^{p}+\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\Big{)}$
$\qquad{}+c(1+\varepsilon^{1-p})\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},\quad\forall\varepsilon\in(0,1).$
Invoking the initial condition $Z_{0}=0$ and exploiting the continuous embedding of $\dot{H}_{h}^{1,q}$ into $L^{q}$, we deduce
$\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\right]\leqslant c\tau\sum_{j=1}^{J-1}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}+c\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$
Applying Proposition 3.1, we infer that
$\tau\sum_{j=1}^{J-1}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\leqslant c\lVert{P_{h}g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$
Moreover, using inequality (3.4) with parameters $\alpha=\beta=0$, we establish
$\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}\leqslant c\lVert{P_{h}g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$
Aggregating these estimates and invoking Lemma 3.1, we conclude
$\mathbb{E}\Big{[}\max_{1\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},$
which implies the stability estimate (3.8). This completes the proof. ∎
3.3 Proof of Theorem 2.1
In this subsection, we always assume that the conditions of Theorem 2.1 are satisfied. Let $y_{h}$ be the solution of the following spatial semi-discretization:
$\begin{cases}{}\mathrm{d}y_{h}(t)=(\Delta_{h}y+y_{h}-P_{h}y_{h}^{3})(t)\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad t>0,\\ y_{h}(0)=P_{h}v.\end{cases}$ (3.14)
It is standard that
$y_{h}(t)$ $=e^{t\Delta_{h}}P_{h}v+\int_{0}^{t}e^{(t-s)\Delta_{h}}(y_{h}-P_{h}y_{h}^{3})(s)\,\mathrm{d}s+{}$ (3.15)
$\qquad\int_{0}^{t}e^{(t-s)\Delta_{h}}P_{h}F(y_{h}(s))\,\mathrm{d}W_{h}(s),\quad t\in[0,T].$
Moreover, Li and Zhou [22, Theorem 3.3] have established the error estimate
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\varepsilon}$ (3.16)
for all $p\in(2,\infty)$, $q\in[2,\infty)$ and $\varepsilon>0$. Hence, to conclude the proof of Theorem 2.1, it remains to prove
$\Big{\lVert}{\max_{1\leqslant j\leqslant J}\lVert{y_{h}(t_{j})-Y_{j}}\rVert_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}\leqslant c\tau^{1/2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty),$ (3.17)
which is the main task of the rest of this subsection.
To this end, we first note that the solution $y$ possesses the following regularity, as indicated by [22, Equation (4.42)]:
$y\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0.$
Using this regularity result, along with the stability property of $P_{h}$ as stated in Lemma 3.1, the standard approximation properties of $P_{h}$, the error estimate Eq. 3.16, and the inverse estimate
$\lVert{v_{h}}\rVert_{\dot{H}_{h}^{1,q}}\leqslant ch^{-1}\lVert{v_{h}}\rVert_{L^{q}},\quad\forall v_{h}\in X_{h},\,\forall q\in[2,\infty),$
we deduce the following regularity assertion for $y_{h}$:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))}+\lVert{P_{h}y_{h}^{3}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))}$ (3.18)
[…]
Nevertheless, it is important to acknowledge that the norm of $y_{h}$ in the space $L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))$ depends on the regularity parameters of the spatial triangulation $\mathcal{K}_{h}$.
The structure of the remaining proof is outlined as follows:
  1. 1.
    In Subsection 3.3.1 we derive the following estimates:
    $\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-y_{h}(t_{j})}\rVert_{L^{q}}^{p}\,\mathrm{d}t\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty),$ (3.19)
    $\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{L^{q}}^{p}\,\mathrm{d}t\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty),$ (3.20)
    $\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1,q}}^{p}\,\mathrm{d}t\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty),$ (3.21)
    $\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\rVert_{L^{q}}^{p}\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty).$ (3.22)
  2. 2.
    In Subsection 3.3.2, we derive the error estimate
    $\max_{0\leqslant j\leqslant J}\lVert{y_{h}(t_{j})-Y_{j}}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\tau^{1/2},\quad\forall p\in[2,\infty).$ (3.23)
  3. 3.
    In Subsection 3.3.3, we establish the stability estimate
    $\hskip-25.00003pt\tau\sum_{j=1}^{J}\mathbb{E}\lVert{Y_{j}}\rVert_{L^{q}}^{p}\text{ is uniformly bounded w.r.t.~{}$h$ and $\tau$ for all $p,q\in[2,\infty)$}.$ (3.24)
  4. 4.
    In Subsection 3.3.4, we establish the error estimate
    $\tau\sum_{j=0}^{J}\mathbb{E}\lVert{Y_{j}-y_{h}(t_{j})}\rVert_{L^{q}}^{p}\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty).$ (3.25)
  5. 5.
    Finally, in Subsection 3.3.5, by utilizing Eqs. 3.18, 3.19, 3.20, 3.21, 3.22, 3.23, 3.24 and 3.25, along with the discrete stochastic maximal $L^{p}$-regularity estimate from Proposition 3.1 and the stability estimate from Proposition 3.2, we conclude the proof for the desired error estimate (3.17).
3.3.1 Proof of the estimates Eqs. 3.19, 3.20, 3.21 and 3.22
In light of the regularity assertion Eq. 3.18 and the inequalities Eqs. 3.1 and 3.2, the derivation of the estimates Eqs. 3.19, 3.20, 3.21 and 3.22 proceeds in a manner that is both direct and routine. Consequently, to maintain brevity, we shall hereinafter provide only a concise proof specifically for the case of Eq. 3.21.
To this end, we proceed as follows. Fix any $0\leqslant j<J$ and $t\in[t_{j},t_{j+1}]$. From Eq. 3.15 we deduce the decomposition
$y_{h}(t_{j+1})-y_{h}(t)=I_{0}+I_{1}+I_{2}+I_{3}+I_{4},$
where each term is defined as follows:
$\hskip-125.00018ptI_{0}:=\Big{(}e^{t_{j+1}\Delta_{h}}-e^{t\Delta_{h}}\Big{)}P_{h}v,$
$\hskip-45.00006ptI_{1}:=\int_{t}^{t_{j+1}}e^{(t_{j+1}-s)\Delta_{h}}\big{(}y_{h}(s)-P_{h}y_{h}^{3}(s)\big{)}\,\mathrm{d}s,$
$I_{2}:=\int_{0}^{t}\Big{(}e^{(t_{j+1}-s)\Delta_{h}}-e^{(t-s)\Delta_{h}}\Big{)}\big{(}y_{h}(s)-P_{h}y_{h}^{3}(s)\big{)}\,\mathrm{d}s,$
$\hskip-50.00008ptI_{3}:=\int_{t}^{t_{j+1}}e^{(t_{j+1}-s)\Delta_{h}}P_{h}F(y_{h}(s))\,\mathrm{d}W_{H}(s),$
$\hskip-5.0ptI_{4}:=\int_{0}^{t}\Big{(}e^{(t_{j+1}-s)\Delta_{h}}-e^{(t-s)\Delta_{h}}\Big{)}P_{h}F(y_{h}(s))\,\mathrm{d}W_{H}(s).$
Using the inequalities Eq. 3.1 and Eq. 3.2, we obtain the following bounds for the terms in the decomposition:
$\lVert{I_{0}}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}$ $\leqslant c\tau^{1/2}\lVert{P_{h}v}\rVert_{\dot{H}_{h}^{2,q}},$
$\lVert{I_{1}}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}$ $\leqslant c\tau\lVert{y_{h}-P_{h}y_{h}^{3}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))},$
$\lVert{I_{2}}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}$ $\leqslant c\tau^{1/2}\lVert{y_{h}-P_{h}y_{h}^{3}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))},$
$\lVert{I_{3}}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}$ $\leqslant c\tau^{1/2}\lVert{P_{h}F(y_{h})}\rVert_{L^{p}(\Omega;C([0,T];\gamma(H,\dot{H}_{h}^{1,q}))},$
$\lVert{I_{4}}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}$ $\leqslant c\tau^{1/2}\lVert{P_{h}F(y_{h})}\rVert_{L^{p}(\Omega;C([0,T];\gamma(H,\dot{H}_{h}^{1,q}))}.$
Drawing upon the stability property of the projection $P_{h}$ as established in Lemma 3.1, the initial condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, the growth estimate Eq. 3.6, and the regularity assertion Eq. 3.18, we demonstrate that the sum of the norms
$\lVert{P_{h}v}\rVert_{\dot{H}_{h}^{2,q}}+\lVert{y_{h}-P_{h}y_{h}^{3}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))}+\lVert{P_{h}F(y_{h})}\rVert_{L^{p}(\Omega;C([0,T];\gamma(H,\dot{H}_{h}^{1,q})))}$
remains uniformly bounded with respect to the mesh size $h$. Consequently, we combine these results to conclude that
$\lVert{y_{h}(t_{j+1})-y_{h}(t)}\rVert_{L^{p}(\Omega;\dot{H}_{h}^{1,q})}\leqslant c\tau^{1/2}$
for all $t\in[t_{j},t_{j+1}]$ with $0\leqslant j<J$. Subsequently, it follows that
$\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1,q}}^{p}\,\mathrm{d}t\leqslant c\sum_{j=0}^{J-1}\tau^{1+p/2}\leqslant c\tau^{p/2},$
thereby establishing the desired estimate Eq. 3.21.
3.3.2 Proof of the error estimate Eq. 3.23
Let $e_{j}:=Y_{j}-y_{h}(t_{j})$ for each $0\leqslant j\leqslant J$. Given Eq. 3.14, we have
$y_{h}(t_{j+1})-y_{h}(t_{j})$ $=\int_{t_{j}}^{t_{j+1}}\Delta_{h}y_{h}(t)+y_{h}(t)-P_{h}y_{h}^{3}(t)\,\mathrm{d}t+P_{h}\int_{t_{j}}^{t_{j+1}}F(y_{h}(t))\,\mathrm{d}W_{H}(t),$
from which, by Eq. 2.2, we deduce that
$e_{j+1}-e_{j}$ $=\int_{t_{j}}^{t_{j+1}}\Delta_{h}\big{(}Y_{j+1}-y_{h}(t)\big{)}+Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\,\mathrm{d}t$ (3.26)
$\quad{}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad\forall 0\leqslant j<J.$
The remainder of the proof is organized into the following three steps.
Step 1. Let us prove
$\mathbb{E}\lVert{e_{j+1}}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\mathbb{E}\big{\lVert}{e_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\big{\rVert}_{L^{2}}^{2m}$ (3.27)
$\qquad{}+c\tau\mathbb{E}\lVert{e_{j}}\rVert_{L^{2}}^{2m}+c\mathbb{E}\big{(}A_{j,m}+B_{j,m}+C_{j,m}\big{)},\quad\forall m\in\mathbb{N}_{>0},\,\,\forall 0\leqslant j<J,$
where
$A_{j,m}$ $:=\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j})}\rVert_{L^{2}}^{2m}\,\mathrm{d}t,$
$B_{j,m}$ $:=\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1,2}}^{2m}\,\mathrm{d}t,$
$C_{j,m}$ $:=\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\rVert_{L^{2}}^{2m}\,\mathrm{d}t.$
Fix any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$. Consider the function
$g(\theta)$ $:=e_{j}+\theta\int_{t_{j}}^{t_{j+1}}\Big{[}\Delta_{h}\big{(}Y_{j+1}-y_{h}(t)\big{)}+Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\Big{]}\,\mathrm{d}t$
$\qquad{}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad\theta\in[0,1].$
For any $\theta\in[0,1]$, differentiating $\lVert{g(\theta)}\rVert_{L^{2}}^{2m}$ with respect to $\theta$ yields
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}$ $=2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\langle{g(\theta),g^{\prime}(\theta)}\rangle$
$=2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\langle{g(1)-(1-\theta)g^{\prime}(\theta),g^{\prime}(\theta)}\rangle$
$\leqslant 2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\langle{g(1),g^{\prime}(\theta)}\rangle,$
where the second equality uses the linearity of $g(\theta)$. Hence, invoking Eq. 3.26 which implies $g(1)=e_{j+1}$, and using the definition of $g(\theta)$, we establish for any $\theta\in[0,1]$ that
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}\leqslant 2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\big{(}I_{1}(\theta)+I_{2}(\theta)\big{)},$
where
$I_{1}(\theta)$ $:=\int_{t_{j}}^{t_{j+1}}\langle{e_{j+1},\,y_{h}^{3}(t_{j+1})-Y_{j+1}^{3}}\rangle\,\mathrm{d}t,$
$I_{2}(\theta)$ $:=\int_{t_{j}}^{t_{j+1}}\big{\langle}{e_{j+1},\,\Delta_{h}(Y_{j+1}-y_{h}(t))+Y_{j}-y_{h}(t)+y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\big{\rangle}.$
Given the inequality
$\langle e_{j+1},\,y_{h}^{3}(t_{j+1})-Y_{j+1}^{3}\rangle\leqslant 0,$
we infer that $I_{1}(\theta)\leqslant 0$. Using the definition of $\Delta_{h}$ and recognizing that $e_{j+1}=Y_{j+1}-y_{h}(t_{j+1})$, we derive the following for any $t\in[t_{j},t_{j+1}]$:
$\big{\langle}{e_{j+1},\,\Delta_{h}(Y_{j+1}-y_{h}(t))+Y_{j}-y_{h}(t)+y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\big{\rangle}$
[…]
$\qquad{}+\big{\langle}{e_{j+1},\,Y_{j}-y_{h}(t)+y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\big{\rangle}.$
Employing Hölder’s inequality, Young’s inequality, and Poincaré’s inequality, we further deduce that
$\big{\langle}{e_{j+1},\,\Delta_{h}(Y_{j+1}-y_{h}(t))+Y_{j}-y_{h}(t)+y_{h}^{3}(t)-Y_{j+1}^{3}}\big{\rangle}$
[…]
Subsequently, we establish that
$I_{2}(\theta)$ $\leqslant c\int_{t_{j}}^{t_{j+1}}\Big{[}\|y_{h}(t)-y_{h}(t_{j+1})\|_{\dot{H}_{h}^{1,2}}^{2}$
$\qquad\qquad+\|Y_{j}-y_{h}(t)\|_{L^{2}}^{2}+\|y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})\|_{L^{2}}^{2}\Big{]}\,\mathrm{d}t.$
By combining the above bounds for $I_{1}$ and $I_{2}$, we derive the differential inequality for $\|g(\theta)\|_{L^{2}}^{2m}$:
$\frac{\mathrm{d}}{\mathrm{d}\theta}\|g(\theta)\|_{L^{2}}^{2m}$ $\leqslant c\|g(\theta)\|_{L^{2}}^{2m-2}\int_{t_{j}}^{t_{j+1}}\Big{[}\|y_{h}(t)-y_{h}(t_{j+1})\|_{\dot{H}_{h}^{1,2}}^{2}$
$\qquad\qquad+\|Y_{j}-y_{h}(t)\|_{L^{2}}^{2}+\|y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})\|_{L^{2}}^{2}\Big{]}\,\mathrm{d}t.$
Applying Young’s inequality and using the definition $e_{j}=Y_{j}-y_{h}(t_{j})$, along with the specified $A_{j,m}$, $B_{j,m}$, and $C_{j,m}$, a straightforward computation yields the differential inequality:
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}$ $\leqslant c\tau\lVert{g(\theta)}\rVert_{L^{2}}^{2m}+c\tau\lVert{e_{j}}\rVert_{L^{2}}^{2m}+c\big{(}A_{j,m}+B_{j,m}+C_{j,m}\big{)}.$
Upon integrating this differential inequality over $[0,s]$ for any $s\in[0,1]$, we attain
$\lVert{g(s)}\rVert_{L^{2}}^{2m}\leqslant\lVert{g(0)}\rVert_{L^{2}}^{2m}+c\int_{0}^{s}\left(\tau\lVert{g(\theta)}\rVert_{L^{2}}^{2m}+\tau\lVert{e_{j}}\rVert_{L^{2}}^{2m}+A_{j,m}+B_{j,m}+C_{j,m}\right)\,\mathrm{d}\theta.$
Employing Gronwall’s lemma subsequently gives
$\lVert{g(1)}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\lVert{g(0)}\rVert_{L^{2}}^{2m}+c\tau\lVert{e_{j}}\rVert_{L^{2}}^{2m}+c\left(A_{j,m}+B_{j,m}+C_{j,m}\right).$
Taking the expectations on both sides yields
$\mathbb{E}\lVert{g(1)}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\mathbb{E}\lVert{g(0)}\rVert_{L^{2}}^{2m}+c\tau\mathbb{E}\lVert{e_{j}}\rVert_{L^{2}}^{2m}+c\mathbb{E}\left(A_{j,m}+B_{j,m}+C_{j,m}\right).$
The desired claim Eq. 3.27 then follows by recognizing that $g(1)=e_{j+1}$ and $g(0)=e_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)$.
Step 2. Fix any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$. For brevity, we denote $P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)$ as $\eta$. With the equality $\mathbb{E}[\lVert{e_{j}}\rVert_{L^{2}}^{2m-2}\big{\langle}{e_{j},\,\eta}\big{\rangle}]=0$, a direct calculation gives
$\mathbb{E}\left\lVert e_{j}+\eta\right\rVert_{L^{2}}^{2m}$ $=\mathbb{E}\left[\left\lVert e_{j}\right\rVert_{L^{2}}^{2}+2\langle e_{j},\eta\rangle+\left\lVert\eta\right\rVert_{L^{2}}^{2}\right]^{m}$
$=\mathbb{E}\left\lVert e_{j}\right\rVert_{L^{2}}^{2m}+\mathbb{E}\Big{[}\sum_{(\alpha,\beta,\gamma)\in\Lambda}C_{\alpha,\beta,\gamma}\left\lVert e_{j}\right\rVert_{L^{2}}^{2\alpha}\langle e_{j},\eta\rangle^{\beta}\left\lVert\eta\right\rVert_{L^{2}}^{2\gamma}\Big{]},$
where $\Lambda$ is the set of all triplets $(\alpha,\beta,\gamma)\in\mathbb{N}^{3}$ satisfying $\alpha+\beta+\gamma=m$ and $\alpha\neq m$, excluding the specific case $(m-1,1,0)$. Each $C_{\alpha,\beta,\gamma}$ is a positive integer constant dependent solely on the triple $(\alpha,\beta,\gamma)$ for each $(\alpha,\beta,\gamma)\in\Lambda$. Applying Cauchy’s inequality followed by Young’s inequality, and considering the constraint $-\frac{2\alpha+\beta}{\beta+2\gamma}\geqslant 1-m$ for all $(\alpha,\beta,\gamma)\in\Lambda$, we deduce that
$\mathbb{E}\bigg{[}\sum_{(\alpha,\beta,\gamma)\in\Lambda}C_{\alpha,\beta,\gamma}\lVert{e_{j}}\rVert_{L^{2}}^{2\alpha}\langle{e_{j},\eta}\rangle^{\beta}\lVert{\eta}\rVert_{L^{2}}^{2\gamma}\bigg{]}$ $\leqslant\mathbb{E}\bigg{[}\sum_{(\alpha,\beta,\gamma)\in\Lambda}C_{\alpha,\beta,\gamma}\lVert{e_{j}}\rVert_{L^{2}}^{2\alpha+\beta}\lVert{\eta}\rVert_{L^{2}}^{\beta+2\gamma}\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\sum_{(\alpha,\beta,\gamma)\in\Lambda}\Big{(}\tau\lVert{e_{j}}\rVert_{L^{2}}^{2m}+\tau^{-\frac{2\alpha+\beta}{\beta+2\gamma}}\lVert{\eta}\rVert_{L^{2}}^{2m}\Big{)}\bigg{]}$
$\leqslant c\tau\mathbb{E}\Big{[}\lVert{e_{j}}\rVert_{L^{2}}^{2m}+\tau^{1-m}\lVert{\eta}\rVert_{L^{2}}^{2m}\Big{]}.$
Consequently, we arrive at the inequality
$\mathbb{E}\big{\lVert}{e_{j}+\eta}\big{\rVert}_{L^{2}}^{2m}\leqslant(1+c\tau)\mathbb{E}\lVert{e_{j}}\rVert_{L^{2}}^{2m}+c\tau^{1-m}\mathbb{E}\lVert{\eta}\rVert_{L^{2}}^{2m}.$
Considering the second term on the right-hand side of this inequality, we infer that
$\tau^{1-m}\mathbb{E}\lVert{\eta}\rVert_{L^{2}}^{2m}$ $\leqslant c\tau^{1-m}\mathbb{E}\Big{[}\int_{t_{j}}^{t_{j+1}}\lVert{F(Y_{j})-F(y_{h}(t))}\rVert_{\gamma(H,L^{2})}^{2}\,\mathrm{d}t\Big{]}^{m}$
$\leqslant c\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{F(Y_{j})-F(y_{h}(t))}\rVert_{\gamma(H,L^{2})}^{2m}\,\mathrm{d}t$
$\leqslant c\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{Y_{j}-y_{h}(t)}\rVert_{L^{2}}^{2m}\,\mathrm{d}t$
$\leqslant c\tau\mathbb{E}\lVert{e_{j}}\rVert_{L^{2}}^{2m}+A_{j,m},$
where the initial inequality is justified by Lemmas 3.1 and 3.3; the subsequent inequality holds due to Hölder’s inequality; the next one follows from Eq. 3.7; and the final inequality stems from the definition of $A_{j,m}$ in conjunction with the observation that $e_{j}=Y_{j}-y_{h}(t_{j})$. Combining the above estimates leads to the final inequality
$\mathbb{E}\big{\lVert}{e_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\big{\rVert}_{L^{2}}^{2m}$ (3.28)
[…]
Step 3. By combining Eq. 3.27 and Eq. 3.28, we obtain that for any positive integer $m$ and for all $0\leqslant j<J$,
$\|e_{j+1}\|_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant(1+c\tau)\|e_{j}\|_{L^{2m}(\Omega;L^{2})}^{2m}+c\mathbb{E}(A_{j,m}+B_{j,m}+C_{j,m}).$
Applying the discrete Gronwall’s lemma and recalling that $e_{0}=0$, we deduce
$\max_{0\leqslant j\leqslant J}\|e_{j}\|_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant c\sum_{j=0}^{J-1}\mathbb{E}(A_{j,m}+B_{j,m}+C_{j,m}),\qquad\forall m\in\mathbb{N}_{>0}.$
Taking into account Eqs. 3.19, 3.21 and 3.22, this establishes the validity of Eq. 3.23 for all $p=2m$, where $m$ is a positive integer. The generalization to the broader case $p\in[2,\infty)$ is achieved by invoking Hölder’s inequality. This completes the proof of the error estimate Eq. 3.23.
3.3.3 Proof of the stability estimate Eq. 3.24
From Eq. 2.2, we decompose the numerical solution into three components:
$Y_{j}=\eta_{j}^{(0)}+\eta_{j}^{(1)}+\eta_{j}^{(2)}\quad\text{for all }1\leqslant j\leqslant J.$
These components are defined by the recursive relations:
$\left\{\begin{array}[]{ll}\eta^{(0)}_{j+1}-\eta^{(0)}_{j}=\tau\Delta_{h}\eta^{(0)}_{j+1},&\quad 0\leqslant j<J,\\ \eta_{0}^{(0)}=P_{h}v,\end{array}\right.$
$\left\{\begin{array}[]{ll}\eta^{(1)}_{j+1}-\eta^{(1)}_{j}=\tau\Delta_{h}\left(\eta^{(1)}_{j+1}+Y_{j}-P_{h}Y_{j+1}^{3}\right),&\quad 0\leqslant j<J,\\ \eta_{0}^{(1)}=0,\end{array}\right.$
and
$\left\{\begin{array}[]{ll}\eta^{(2)}_{j+1}-\eta^{(2)}_{j}=\tau\Delta_{h}\eta^{(2)}_{j+1}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})\,\mathrm{d}W_{H}(t),&\quad 0\leqslant j<J,\\ \eta_{0}^{(2)}=0.\end{array}\right.$
Given the identity $\eta_{j}^{(0)}=(I-\tau\Delta_{h})^{-j}P_{h}v$ holding for all $0\leqslant j\leqslant J$, and using the stability property of $P_{h}$ from Lemma 3.1 along with Eq. 3.4 with $\alpha=\beta=0$, we assert that
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\eta_{j}^{(0)}}\rVert_{L^{q}}^{p}\text{ is uniformly bounded w.r.t.~{}$h$ and $\tau$ for all $p,q\in[2,\infty)$}.$ (3.29)
For any $p\in[2,\infty)$, from Lemma 3.1 and Eq. 3.5, it directly follows that
$\tau\sum_{j=0}^{J}\mathbb{E}\left[\lVert{Y_{j}-P_{h}Y_{j}^{3}}\rVert_{L^{2}}^{p}+\lVert{P_{h}F(Y_{j})}\rVert_{\gamma(H,L^{6})}^{p}\right]$ $\leqslant c\tau\sum_{j=0}^{J}\mathbb{E}\left[1+\lVert{Y_{j}}\rVert_{L^{6}}^{3p}\right].$
Moreover, one can establish that
$\tau\sum_{j=0}^{J}\mathbb{E}\lVert{Y_{j}}\rVert_{L^{6}}^{3p}$ $\leqslant c\max_{0\leqslant j\leqslant J}\lVert{Y_{j}}\rVert_{L^{3p}(\Omega;L^{6})}^{3p}$
$\leqslant c\max_{0\leqslant j\leqslant J}\left[\lVert{Y_{j}-y_{h}(t_{j})}\rVert_{L^{3p}(\Omega;L^{6})}^{3p}+\lVert{y_{h}(t_{j})}\rVert_{L^{3p}(\Omega;L^{6})}^{3p}\right]$
$\leqslant c\max_{0\leqslant j\leqslant J}\left[1+\lVert{y_{h}(t_{j})}\rVert_{L^{3p}(\Omega;L^{6})}^{3p}\right]$
$\leqslant c\left(1+\lVert{y_{h}}\rVert_{L^{3p}(\Omega;C([0,T];L^{6}))}^{3p}\right),$
where the third inequality is justified by Eq. 3.23, the condition $\tau\leqslant h^{2}$, and the application of the inverse inequality $\lVert{v_{h}}\rVert_{L^{6}}\leqslant ch^{-1}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$. Consequently, by appealing to the regularity assertion in Eq. 3.18, we conclude that
$\tau\sum_{j=0}^{J}\mathbb{E}\left[\lVert{Y_{j}-P_{h}Y_{j}^{3}}\rVert_{L^{2}}^{p}+\lVert{P_{h}F(Y_{j})}\rVert_{\gamma(H,L^{6})}^{p}\right]$
remains uniformly bounded with respect to $h$ and $\tau$ for all $p\in[2,\infty)$. Using this result in conjunction with Lemmas 3.5 and 3.1, and taking advantage of the continuous embeddings of $\dot{H}_{h}^{2,2}$ and $\dot{H}_{h}^{1,6}$ into $L^{q}$ for any $q\in[2,\infty)$, we are able to assert that
$\tau\sum_{j=1}^{J}\mathbb{E}\left[\|\eta_{j}^{(1)}\|_{L^{q}}^{p}+\|\eta_{j}^{(2)}\|_{L^{q}}^{p}\right]$
is also uniformly bounded with respect to $h$ and $\tau$ for all $p,q\in[2,\infty)$. This finding, when coupled with the assertion in Eq. 3.29 and recognizing that $Y_{j}=\eta_{j}^{(0)}+\eta_{j}^{(1)}+\eta_{j}^{(2)}$ for $1\leqslant j\leqslant J$, permits us to affirm the stability estimate stated in Eq. 3.24.
3.3.4 Proof of the error estimate Eq. 3.25
Firstly, according to Eq. 3.26, for each $0\leqslant j\leqslant J$, we decompose the error $Y_{j}-y_{h}(t_{j})$ into three components:
$Y_{j}-y_{h}(t_{j})=\xi_{j}^{(0)}+\xi_{j}^{(1)}+\xi_{j}^{(2)}.$ (3.30)
The sequences $(\xi_{j}^{(0)})_{j=0}^{J}$, $(\xi_{j}^{(1)})_{j=0}^{J}$, and $(\xi_{j}^{(2)})_{j=0}^{J}$ are recursively defined as follows:
$\begin{cases}\xi^{(0)}_{j+1}-\xi^{(0)}_{j}=\tau\Delta_{h}\xi^{(0)}_{j+1}+\int_{t_{j}}^{t_{j+1}}\Delta_{h}\big{(}y_{h}(t_{j+1})-y_{h}(t)\big{)}\,\mathrm{d}t,\quad 0\leqslant j<J,\\ \xi_{0}^{(0)}=0,\end{cases}$
[…]
and
$\begin{cases}\xi^{(2)}_{j+1}-\xi^{(2)}_{j}=\tau\Delta_{h}\xi^{(2)}_{j+1}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant j<J,\\ \xi_{0}^{(2)}=0.\end{cases}$
Secondly, by invoking Lemma 3.5 and using the assertion presented in Eq. 3.20, we derive the following bound for $(\xi_{j}^{(0)})_{j=1}^{J}$:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(0)}}\rVert_{L^{q}}^{p}$ $\leqslant\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{L^{q}}^{p}\,\mathrm{d}t$
$\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty).$ (3.31)
Thirdly, we proceed to estimate the sequence $(\xi_{j}^{(1)})_{j=1}^{J}$. Let us fix an $p,q\in[2,\infty)$. A direct computation yields
$\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\,\mathrm{d}t$
[…]
$\qquad\qquad\qquad\quad{}+\lVert{P_{h}y_{h}^{3}(t)-P_{h}y_{h}^{3}(t_{j+1})+P_{h}y_{h}^{3}(t_{j+1})-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\bigg{]}\,\mathrm{d}t$
[…]
$\qquad\qquad\qquad\quad{}+\lVert{P_{h}y_{h}^{3}(t)-P_{h}y_{h}^{3}(t_{j+1})}\rVert_{L^{3/2}}^{p}+\lVert{P_{h}y_{h}^{3}(t_{j+1})-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\bigg{]}\,\mathrm{d}t.$
Invoking the stability property of $P_{h}$ as established in Lemma 3.1, and utilizing the results from Eq. 3.19, Eq. 3.22, and Eq. 3.23, we further infer that
$\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\,\mathrm{d}t$
[…]
Applying Hölder’s inequality to the cubic difference term, we obtain the following bound:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{y_{h}^{3}(t_{j})-Y_{j}^{3}}\rVert_{L^{3/2}}^{p}$
[…]
By using the regularity assertion in Eq. 3.18, the stability estimate Eq. 3.24, and the error estimate Eq. 3.23, we establish that
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{y_{h}^{3}(t_{j})-Y_{j}^{3}}\rVert_{L^{3/2}}^{p}\leqslant c\tau^{p/2}.$
Aggregating these inequalities, we arrive at
$\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\,\mathrm{d}t\leqslant c\tau^{p/2}.$ (3.32)
Employing Lemma 3.5 and the continuous embedding of $\dot{H}_{h}^{2,3/2}$ in $L^{q}$, we deduce that
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(1)}}\rVert_{L^{q}}^{p}$ $\leqslant c\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(1)}}\rVert_{\dot{H}_{h}^{2,3/2}}^{p}$
$\leqslant c\sum_{j=0}^{J-1}\mathbb{E}\int_{t_{j}}^{t_{j+1}}\lVert{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{3/2}}^{p}\,\mathrm{d}t.$
Merging this inequality with Eq. 3.32 leads to the bound for $(\xi_{j}^{(1)})_{j=1}^{J}$:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(1)}}\rVert_{L^{q}}^{p}\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty).$ (3.33)
Fourthly, concerning the sequence $(\xi_{j}^{(2)})_{j=1}^{J}$, for any $p\in[2,\infty)$, we derive:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(2)}}\rVert_{L^{6}}^{p}$ $\leqslant c\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(2)}}\rVert_{\dot{H}_{h}^{1,2}}^{p}$
$\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{P_{h}\big{(}F(Y_{j})-F(y_{h}(t))\big{)}}\rVert_{\gamma(H,L^{2})}^{p}\,\mathrm{d}t$
$\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{y_{h}(t)-Y_{j}}\rVert_{L^{2}}^{p}\,\mathrm{d}t,$
where the first inequality uses the continuous embedding of $\dot{H}_{h}^{1,2}$ into $L^{6}$, the second inequality follows from Proposition 3.1, and the third inequality is a direct consequence of Lemmas 3.1 and 3.7. Subsequently, by Eqs. 3.19 and 3.23 we establish the following bound for $(\xi_{j}^{(2)})_{j=1}^{J}$:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(2)}}\rVert_{L^{6}}^{p}\leqslant c\tau^{p/2},\quad\forall p\in[2,\infty).$
Combining this bound with Eqs. 3.30, 3.31 and 3.33 yields the error estimate
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{Y_{j}-y_{h}(t_{j})}\rVert_{L^{6}}^{p}\leqslant c\tau^{p/2},\quad\forall p\in[2,\infty).$
Employing this estimate and following a reasoning process analogous to that used in the derivation of the aforementioned bound for $(\xi_{j}^{(2)})_{j=1}^{J}$, we obtain a refined bound:
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{\xi_{j}^{(2)}}\rVert_{L^{q}}^{p}\leqslant c\tau^{p/2},\quad\forall p,q\in[2,\infty).$ (3.34)
Finally, in light of the decomposition Eq. 3.30 and the identity $Y_{0}=y_{h}(0)$, the desired error estimate Eq. 3.25 is readily established by synthesizing the bounds established in Eqs. 3.31, 3.33 and 3.34.
3.3.5 Proof of the error estimate Eq. 3.17
Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. Let the sequences $(\xi_{j}^{(0)})_{j=0}^{J}$, $(\xi_{j}^{(1)})_{j=0}^{J}$, and $(\xi_{j}^{(2)})_{j=0}^{J}$ be constructed as in Subsection 3.3.4. By their construction, for any $1\leqslant j\leqslant J$, the $L^{q}$-norm of $\xi_{j}^{(0)}$ can be estimated as follows:
$\lVert{\xi_{j}^{(0)}}\rVert_{L^{q}}$ $\leqslant\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{(I-\tau\Delta_{h})^{k-j}\Delta_{h}\left(y_{h}(t_{k+1})-y_{h}(t)\right)}\rVert_{L^{q}}\,\mathrm{d}t$
$=\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{(I-\tau\Delta_{h})^{k-j}\left(y_{h}(t_{k+1})-y_{h}(t)\right)}\rVert_{\dot{H}_{h}^{2,q}}\,\mathrm{d}t$
$\leqslant\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{(I-\tau\Delta_{h})^{k-j}}\rVert_{\mathcal{L}(\dot{H}_{h}^{1,q},\dot{H}_{h}^{2,q})}\lVert{y_{h}(t)-y_{h}(t_{k+1})}\rVert_{\dot{H}_{h}^{1,q}}\,\mathrm{d}t.$
Applying the resolvent estimate Eq. 3.4 with parameters $\alpha=1$ and $\beta=2$, invoking Hölder’s inequality, and acknowledging the uniform boundedness of the integral
$\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\big{(}(j-k)\tau\big{)}^{-\frac{p}{2(p-1)}}\,\mathrm{d}t$
with respect to the time step $\tau$, we deduce that for any $1\leqslant j\leqslant J$,
$\lVert{\xi_{j}^{(0)}}\rVert_{L^{q}}$ $\leqslant c\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\big{(}(j-k)\tau\big{)}^{-\frac{1}{2}}\lVert{y_{h}(t)-y_{h}(t_{k+1})}\rVert_{\dot{H}_{h}^{1,q}}\,\mathrm{d}t$
$\leqslant c\left[\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{y_{h}(t)-y_{h}(t_{k+1})}\rVert_{\dot{H}_{h}^{1,q}}^{p}\,\mathrm{d}t\right]^{\frac{1}{p}}.$
Thus, for $(\xi_{j}^{(0)})_{j=1}^{J}$, we deduce the bound:
$\max_{1\leqslant j\leqslant J}\lVert{\xi_{j}^{(0)}}\rVert_{L^{q}}^{p}\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1,q}}^{p}\,\mathrm{d}t.$
Analogously, it is established that
$\max_{1\leqslant j\leqslant J}\lVert{\xi_{j}^{(1)}}\rVert_{L^{q}}^{p}\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{q}}^{p}\,\mathrm{d}t.$
Combining these results leads to
$\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\lVert{\xi_{j}^{(0)}+\xi_{j}^{(1)}}\rVert_{L^{q}}^{p}\right]$
[…]
Furthermore, utilizing Eq. 3.25 and adopting the proof technique employed for Eq. 3.32, we establish the following inequality
$\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\big{\lVert}{Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\big{\rVert}_{L^{q}}^{p}\,\mathrm{d}t\leqslant c\tau^{p/2}.$
Hence, we conclude that
$\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\big{\lVert}{\xi_{j}^{(0)}+\xi_{j}^{(1)}}\big{\rVert}_{L^{q}}^{p}\right]\leqslant c\tau^{p/2},$
which implies the bound
$\Big{\lVert}{\max_{1\leqslant j\leqslant J}\big{\lVert}{\xi_{j}^{(0)}+\xi_{j}^{(1)}}\big{\rVert}_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}\leqslant c\tau^{1/2}.$ (3.35)
Next, regarding the sequence $(\xi_{j}^{(2)})_{j=1}^{J}$, by invoking the stability estimate established in Lemma 3.2, we deduce the following bound:
$\left\|\max_{1\leqslant j\leqslant J}\|\xi_{j}^{(2)}\|_{L^{q}}\right\|_{L^{p}(\Omega)}$ $\leqslant c\bigg{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\left\|P_{h}\big{(}F(y_{h}(t))-F(Y_{j})\big{)}\right\|_{\gamma(H,L^{q})}^{p}\,\mathrm{d}t\bigg{]}^{1/p}.$
Using the stability property of $P_{h}$ from Lemma 3.1 and the Lipschitz continuity of $F$ given by Eq. 3.7, we proceed to refine this bound:
$\Big{\|}\max_{1\leqslant j\leqslant J}\|\xi_{j}^{(2)}\|_{L^{q}}\Big{\|}_{L^{p}(\Omega)}$ $\leqslant c\bigg{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\left\|y_{h}(t)-Y_{j}\right\|_{L^{q}}^{p}\,\mathrm{d}t\bigg{]}^{1/p}$
$\leqslant c\bigg{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\left\|y_{h}(t)-y_{h}(t_{j})\right\|_{L^{q}}^{p}\,\mathrm{d}t\bigg{]}^{1/p}$
$\phantom{{}\leqslant{}}+c\bigg{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\mathbb{E}\left\|y_{h}(t_{j})-Y_{j}\right\|_{L^{q}}^{p}\,\mathrm{d}t\bigg{]}^{1/p}.$
Furthermore, applying the results from Eq. 3.19 and Eq. 3.25, we establish the bound:
$\Big{\|}\max_{1\leqslant j\leqslant J}\|\xi_{j}^{(2)}\|_{L^{q}}\Big{\|}_{L^{p}(\Omega)}\leqslant c\tau^{1/2}.$ (3.36)
Finally, given that the decomposition Eq. 3.30 implies
$\Big{\lVert}{\max_{1\leqslant j\leqslant J}\big{\lVert}{Y_{j}-y_{h}(t_{j})}\big{\rVert}_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}$
[…]
the desired error estimate Eq. 3.17 is readily obtained by combining the bounds Eq. 3.35 and Eq. 3.36. This completes the proof of Theorem 2.1.
4 Numerical results
This section is dedicated to the numerical validation of the theoretical findings presented in this paper. The numerical experiments are conducted using the following model equation:
$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+y(t)\,\mathrm{d}\beta(t),\quad 0<t\leqslant T,\\ y(0)=v.\end{cases}$
Here, $\beta$ represents a given Brownian motion, the spatial domain $\mathcal{O}$ is $(0,1)^{3}$, and the initial condition is specified by the function $v$ defined as:
$v(x):=\sin(\pi x_{1})\sin(\pi x_{2})\sin(\pi x_{3})\quad\text{for all }x:=(x_{1},x_{2},x_{3})\in\mathcal{O}.$
To verify the spatial accuracy as established in Theorem 2.1, we set the final time to $T=0.01$ and the time step to $\tau=1\times 10^{-6}$. We then use the numerical solution computed with a spatial mesh size of $h=\frac{1}{2^{8}}$ as our reference solution. Furthermore, we perform simulations across $500$ independent paths. The numerical results presented in Figure 1 agrees well with the expected spatial convergence rate of $O(h^{2-\epsilon})$, as outlined in Eq. 2.3, for the cases of $p=q=2$, $p=q=4$, and $p=q=16$.
To further affirm the temporal accuracy claimed in Theorem 2.1, we set the terminal time to $T=0.1$ and adjust the spatial mesh size $h$ to meet the condition $h=\tau^{1/2}$, ensuring that the temporal discretization error is predominant. We use the numerical solution with a time step of $\tau=T/2^{13}$ as the reference solution. As before, we conduct simulations over $500$ independent paths. The numerical results displayed in Figure 2 confirm the temporal convergence rate of $O(\tau^{1/2})$, consistent with the theoretical convergence rate specified in Eq. 2.3.
Figure 1: $\tau=1.0\times 10^{-6}$, $T=0.01$,
Figure 2: $T=0.1$, $h=\tau^{1/2}$
5 Conclusions
In this paper, a novel stability estimate for a discrete stochastic convolution has been derived as an independent contribution. This stability estimate plays a crucial role in establishing pathwise uniform convergence estimates for fully discrete approximations of nonlinear stochastic parabolic equations. By combining this stability estimate with the discrete stochastic maximal $L^{p}$ regularity estimate, a pathwise uniform convergence rate of $O\big{(}h^{2-\epsilon}+\tau^{1/2}\big{)}$ has been achieved for a full discretization of a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. Moreover, the convergence estimate employs general spatial $L^{q}$-norms. This theoretical finding not only advances our understanding of numerical methods for the stochastic Allen-Cahn equation but also introduces a novel approach for the numerical analysis of nonlinear stochastic parabolic equations, using the discrete stochastic maximal $L^{p}$ regularity estimate.
Funding
The first author was partially supported by Natural Science Foundation of Sichuan Province under grant 2023NSFSC1324. The second author was partially supported by National Natural Science Foundation of China under grant 12301525.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Declaration
Conflict of interest The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.