### v5

# Pathwise uniform convergence of a full discretization for a three-dimensional stochastic Allen-Cahn equation with multiplicative noise

Abstract
This paper analyzes a full discretization of a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. The discretization combines the Euler scheme for temporal approximation and the finite element method for spatial approximation. A pathwise uniform convergence rate is derived, encompassing general spatial $L^{q}$-norms, by using discrete versions of deterministic and stochastic maximal $L^{p}$-regularity estimates. Additionally, the theoretical convergence rate is validated through numerical experiments. The primary contribution of this work is the introduction of a technique to establish the pathwise uniform convergence of finite element-based full discretizations for nonlinear stochastic parabolic equations within the framework of general spatial $L^{q}$-norms.
Keywords: stochastic Allen-Cahn equation, Euler scheme, finite element method, discrete stochastic maximal $L^{p}$-regularity, pathwise uniform convergence
1 Introduction
Let $T\in(0,\infty)$ denote the terminal time, and let $\mathcal{O}$ be a bounded, convex domain in $\mathbb{R}^{3}$ with a smooth boundary $\partial\mathcal{O}$. Let $\Delta$ represent the Laplace operator subject to homogeneous Dirichlet boundary conditions. This paper studies the following stochastic Allen-Cahn equation:
$\begin{cases}\mathrm{d}y(t)=\left(\Delta y+y-y^{3}\right)(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),&t\in[0,T],\\ y(0)=v,\end{cases}$ (1.1)
where $v$ denotes the prescribed initial condition,a and $W_{H}$ represents an $H$-cylindrical Brownian motion. The diffusion coefficient $F$ will be precisely defined later.
The Allen-Cahn equation is an important mathematical model in materials science [1] and it is also of fundamental importance for the geometric moving interface problem [10]. By incorporating thermal fluctuations, impurities, and atomistic processes that describe surface motion, the stochastic Allen-Cahn equation naturally arises [12, 22, 37].
Recently, there has been a growing interest in the numerical analysis of the stochastic Allen-Cahn equation. A brief summary of some recent works is given below. For the stochastic Allen-Cahn equation with additive noise, Becker and Jentzen [4] obtained strong convergence rates for a nonlinearity-truncated Euler-type approximation. Becker et al. [3] established sharp strong convergence rates for an explicit space-time discrete numerical approximation. Bréhier et al. [5] analyzed an explicit temporal splitting numerical scheme. Qi and Wang [35] derived strong error estimates for a full discretization using the finite element method for spatial discretization and the Euler scheme for temporal discretization. In addition, Wang [43] analyzed an efficient explicit full-discrete scheme for a one-dimensional stochastic Allen–Cahn equation. The numerical analysis of the stochastic Allen-Cahn equation with multiplicative noise generally presents greater challenges. In this regard, Feng et al. [11] derived the strong convergence of finite element approximations of a stochastic Allen-Cahn equation with finite-dimensional gradient-type multiplicative noise, Majee and Prohl [33] analyzed a space-time discretization, and Liu and Qiao [32] investigated the strong convergence of Galerkin-based Euler approximations. We also refer the reader to [14, 15, 17, 18, 19, 23, 24, 25, 36, 38] and the references therein for more related works.
So far, the numerical analysis of the stochastic Allen-Cahn equation in the literature has generally considered only the $L^{p}(\Omega;L^{2})$-norm error estimates for fully discrete approximations. There is a lack of pathwise uniform convergence estimates and convergence estimates with general spatial $L^{q}$-norms. Notably, Bréhier et al. [5] derived some pathwise uniform convergence estimates for an explicit temporal splitting numerical scheme; however, their analysis might not be directly applicable to general domains (see [5, pp. 2116]). Extending this analysis to the full discretization of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise is nontrivial.
The main contributions of this paper are summarized as follows:
  • We establish the stability estimate
    $\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\left\|\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}(I-\tau\Delta_{h})^{k-j}g_{h}(t)\,\mathrm{d}W_{H}(t)\right\|_{L^{q}}^{p}\right]\leqslant C\left\|g_{h}\right\|_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},$
    for all $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,X_{h}))$, where $p\in(2,\infty)$ and $q\in[2,\infty)$. Here, $X_{h}$ denotes the finite element space and $L^{q}$ denotes $L^{q}(\mathcal{O})$. This estimate is useful for demonstrating pathwise uniform convergence of fully discrete approximations based on finite elements for nonlinear stochastic parabolic equations, within the framework of general spatial $L^{q}$-norms. While van Neerven and Veraar [40] have developed analogous estimates for the Euler scheme, their methodology does not directly apply to deriving the aforementioned stability estimate, as detailed in Subsection 4.2.
  • For a fully discrete approximation of Eq. 1.1, using the Euler scheme for temporal discretization and the finite element method for spatial discretization, we establish the following pathwise uniform convergence estimate:
    $\Big{\lVert}{\max_{1\leqslant j\leqslant J}\lVert{y(t_{j})-Y_{j}}\rVert_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}\leqslant c\left(h^{2-\varepsilon}+\tau^{1/2}\right)$
    for all $p\in(2,\infty)$, $q\in[2,\infty)$ and $\varepsilon>0$. Notably, even for the three-dimensional stochastic Allen-Cahn equation with additive noise, to the best of our knowledge, such error estimates for fully discrete approximations using the Euler scheme have not been reported previously. This finding fills an important gap in the existing literature.
Our analysis is based on discrete deterministic and stochastic maximal $L^{p}$-regularity estimates. We impose relatively mild regularity conditions on the diffusion coefficient $F$ in Eq. 1.1. Although our analysis assumes that the initial value $v$ is deterministic and belongs to $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, it can be extended to cases where the initial value $v$ has lower regularity. In addition, the pathwise uniform convergence of numerical approximations for SPDEs within the framework of Banach spaces has recently attracted considerable attention; see, for example, [7, 8, 21, 40]. However, existing research has primarily focused on temporal semidiscretizations. This work introduces a new approach for establishing pathwise uniform convergence for finite element-based full discretizations of nonlinear stochastic parabolic equations, within the general spatial $L^{q}$-norms framework.
The remainder of this paper is organized as follows. In Section 2, we introduce the notational conventions used throughout this work and present a regularity result for the model problem. In Section 3, we describe a full discretization and provide a pathwise uniform convergence estimate. In Section 4, we establish a stability estimate for a discrete stochastic convolution and provide a proof for the previously mentioned convergence estimate. Section 5 presents numerical experiments to validate the theoretical result. Finally, Section 6 summarizes the findings and contributions of this study.
2 Mathematical Setup
For any two Banach spaces $E_{1}$ and $E_{2}$, we denote by $\mathcal{L}(E_{1},E_{2})$ the space of all bounded linear operators from $E_{1}$ to $E_{2}$, and abbreviate $\mathcal{L}(E_{1},E_{1})$ to $\mathcal{L}(E_{1})$. The symbol $I$ represents the identity map. Let $\mathcal{O}$ be a bounded, convex domain in $\mathbb{R}^{3}$ of class $\mathcal{C}^{2}$ (cf. Section 11.1 in Chapter 1 of [44]). For any $q\in[1,\infty]$, let $W_{0}^{1,q}(\mathcal{O})$ and $W^{2,q}(\mathcal{O})$ denote the usual Sobolev spaces (see, e.g., [39]), and we shall abbreviate $L^{q}(\mathcal{O})$ to $L^{q}$ for simplicity. Let $\Delta$ denote the Laplace operator subject to homogeneous Dirichlet boundary conditions. For $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the norm
$\lVert{u}\rVert_{\dot{H}^{\alpha,q}}:=\lVert{(-\Delta)^{\alpha/2}u}\rVert_{L^{q}},\quad\forall u\in\dot{H}^{\alpha,q}.$
We use $\dot{H}^{-\alpha,q^{\prime}}$ to denote the dual space of $\dot{H}^{\alpha,q}$, where $q^{\prime}$ is the conjugate exponent of $q$, i.e., $11/q+1/q^{\prime}=$ 1. For any $q\in(1,\infty)$, it is well-known (see, e.g., Theorem 16.14 of [44]) that $\dot{H}^{2,q}$ coincides with the intersection $W_{0}^{1,q}(\mathcal{O})\cap W^{2,q}(\mathcal{O})$, and is equipped with a norm equivalent to that of $W^{2,q}(\mathcal{O})$. Moreover, $\dot{H}^{1,q}$ is identical to $W_{0}^{1,q}(\mathcal{O})$, with equivalent norms.
Assume that $(\Omega,\mathcal{F},\mathbb{P})$ is a given complete probability space with a right-continuous filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ are defined. For any $0\leqslant s<t<\infty$ and for any $n\in\mathbb{N}$, the increment $\beta_{n}(t)-\beta_{n}(s)$ is independent of $\mathcal{F}_{s}$. Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$. Let the $H$-cylindrical Brownian motion be defined such that for each $t\in\mathbb{R}_{+}$, $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ is given by
$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H.$
We use the symbol $\mathbb{E}$ to represent the expectation operator associated with the probability space $(\Omega,\mathcal{F},\mathbb{P})$. For any UMD Banach space $E$, let $\gamma(H,E)$ denote the space of all $\gamma$-radonifying operators from $H$ to $E$. For any $p\in[2,\infty)$, we denote by $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,E))$ and $L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,E)))$ the spaces of all $\mathbb{F}$-adapted $\gamma(H,E)$-valued processes in $L^{p}(\Omega\times(0,T);\gamma(H,E))$ and $L^{p}(\Omega;L^{2}(0,T;\gamma(H,E)))$, respectively. When $E$ is a Hilbert space, $\gamma(H,E)$ coincides with the standard Hilbert-Schmidt space from $H$ to $E$, and we have the celebrated Itô’s isometry:
$\mathbb{E}\left\lVert\int_{0}^{T}g(t)\,\mathrm{d}W_{H}(t)\right\rVert_{E}^{2}=\mathbb{E}\int_{0}^{T}\lVert{g(t)}\rVert_{\gamma(H,E)}^{2}\,\mathrm{d}t,$
for all $g\in L_{\mathbb{F}}^{2}(\Omega\times(0,T);\gamma(H,E))$. For an in-depth discussion of $\gamma$-radonifying operators, the reader is referred to Chapter 9 of [16]. An extensive treatment of stochastic integrals with respect to $W_{H}$ in UMD Banach spaces can be found in [41].
Next, we define the diffusion coefficient $F$ in Eq. 1.1 as follows. Let $(f_{n})_{n\in\mathbb{N}}$ denote a sequence of continuous functions from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$ such that for each $n\in\mathbb{N}$,
$\begin{cases}f_{n}(x,0)=0,&\forall x\in\partial\mathcal{O},\\ \lvert{f_{n}(x,y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert),&\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},\\ \lvert{f_{n}(x,y_{1})-f_{n}(x,y_{2})}\rvert\leqslant C_{F}\lvert{y_{1}-y_{2}}\rvert,&\forall x\in\mathcal{O},\,\forall y_{1},y_{2}\in\mathbb{R},\\ \lvert{f_{n}(x_{1},y)-f_{n}(x_{2},y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert)\lvert{x_{1}-x_{2}}\rvert,&\forall x_{1},x_{2}\in\mathcal{O},\,\forall y\in\mathbb{R},\end{cases}$ (2.1)
where $C_{F}$ is a positive constant independent of $n$, $x$, $y$, $x_{1}$, $x_{2}$, $y_{1}$ and $y_{2}$. Let $(\lambda_{n})_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying the condition $\sum_{n\in\mathbb{N}}\lambda_{n}<\infty$. Finally, for any $u\in L^{2}$ define
$F(u):=\sum_{n=0}^{\infty}\sqrt{\lambda}_{n}h_{n}\otimes f_{n}(\cdot,u(\cdot)),$
where $h_{n}\otimes f_{n}(\cdot,u(\cdot))\in\gamma(H,L^{2})$ is defined by
$\big{(}h_{n}\otimes f_{n}(\cdot,u(\cdot))\big{)}(w):=(w,h_{n})_{H}f_{n}(\cdot,u(\cdot)),\quad\forall w\in H.$
By [16, Proposition 9.1.9] and (2.1), a straightforward calculation gives the following growth property:
$\lVert{F(u)}\rVert_{\gamma(H,L^{2})}\leqslant c\big{(}1+\lVert{u}\rVert_{L^{2}}\big{)},\quad\forall u\in L^{2},$
where $c$ is a positive constant independent of $u$.
We call a process $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ a mild solution to the model problem Eq. 1.1 if it satisfies almost surely the following equality for all $t\in[0,T]$:
$y(t)=e^{t\Delta}v+\int_{0}^{t}e^{(t-s)\Delta}\big{[}y(s)-y^{3}(s)\big{]}\,\mathrm{d}s+\int_{0}^{t}e^{(t-s)\Delta}F(y(s))\,\mathrm{d}W_{H}(s),$
where $e^{t\Delta}$, $t\in[0,T]$, denotes the analytic semigroup generated by the Laplace operator $\Delta$ with homogeneous Dirichlet boundary conditions, and $v$ is the given initial value. While the condition $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ can be relaxed (see, e.g., Chapter 6 of [34]), it suffices for our purposes. Following the proof of Theorem 3.1 in [28], we establish the following regularity result. For related regularity results, we also refer the reader to [32, Section 3] and [30].
Proposition 2.1.
Suppose the initial value $v$ belongs to $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Then, the model problem (1.1) admits a unique mild solution. Moreover, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the solution $y$ satisfies
$y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q})\cap L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})).$
3 Full Discretization
Let $J$ be a positive integer, and set $t_{j}:=j\tau$ for each $0\leqslant j\leqslant J$, where $\tau:=T/J$ is the time step. Let $\mathcal{K}_{h}$ be a conforming and quasi-uniform triangulation of $\mathcal{O}$ consisting of three-dimensional simplexes, and let the spatial mesh size $h$ denote the maximum diameter of the simplexes in $\mathcal{K}_{h}$. Define
$X_{h}:=\Big{\{}u_{h}\in C(\overline{\mathcal{O}})\,\big{|}\,u_{h}=0\text{ on }\partial\mathcal{O}_{h}\text{ and outside of }\mathcal{O}_{h},\text{ and }u_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\Big{\}},$
where $C(\overline{\mathcal{O}})$ is the set of all continuous functions on the closure of $\mathcal{O}$, $\mathcal{O}_{h}$ is the closure of the union of the simplexes in $\mathcal{K}_{h}$, $r\geqslant 1$ is an integer, and $P_{r}(K)$ is the set of all polynomials on $K$ with degree not exceeding $r$ for each $K\in\mathcal{K}_{h}$. Let $P_{h}$ be the $L^{2}$-orthogonal projection operator onto $X_{h}$, and, for any $u_{h}\in X_{h}$, define $\Delta_{h}u_{h}\in X_{h}$ by
$\int_{\mathcal{O}}(\Delta_{h}u_{h})\cdot v_{h}\,\mathrm{d}\mu=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}\mu,\quad\forall v_{h}\in X_{h},$
where $\mu$ denotes the three-dimensional Lebesgue measure on $\mathcal{O}$.
This study considers the following full discretization of the model problem Eq. 1.1:
$\begin{cases}Y_{j+1}-Y_{j}=\tau\left(\Delta_{h}Y_{j+1}+Y_{j}-P_{h}Y_{j+1}^{3}\right)+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})\,\mathrm{d}W_{H}(t),\quad 0\leqslant j<J,\\ Y_{0}=P_{h}v.\end{cases}$ (3.1)
For this full discretization, we establish the following pathwise uniform convergence estimate; the proof is deferred to Section 4.
Theorem 3.1.
Assume that the initial value $v$ belongs to $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, and $\tau\leqslant h^{2}$. Let $y$ be the mild solution of the model problem Eq. 1.1, and let $(Y_{j})_{j=0}^{J}$ be the solution of the full discretization Eq. 3.1. Then for all $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following pathwise uniform convergence estimate holds:
$\Big{\lVert}{\max_{1\leqslant j\leqslant J}\lVert{Y_{j}-y(t_{j})}\rVert_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}\leqslant c\big{(}h^{2-\varepsilon}+\tau^{1/2}\big{)},$ (3.2)
where $c$ denotes a positive constant independent of the spatial mesh size $h$ and the time step $\tau$.
Remark 3.1.
For Theorem 3.1, we have the following comments.
  1. 1.
    The condition $\tau\leqslant h^{2}$ can be relaxed by re-examining and refining the analysis in Subsection 4.3.2.
  2. 2.
    Assume that the sequence $(f_{n})_{n\in\mathbb{N}}$, in addition to satisfying (2.1), consists of twice continuously differentiable functions and satisfies the following growth condition for each $n\in\mathbb{N}$:
    $\left|\nabla_{x}^{2}f_{n}(x,y)\right|+\left|\nabla_{x}\nabla_{y}f_{n}(x,y)\right|+\left|\nabla_{y}^{2}f_{n}(x,y)\right|\leqslant C_{F}(1+|y|),\quad\forall x\in\mathcal{O},\,\forall y\in\mathbb{R}^{3},$
    where $C_{F}$ is a positive constant independent of $n$. Under the condition that the initial value $v$ belongs to $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, the spatial accuracy $O(h^{2-\epsilon})$ in Eq. 3.2 can be improved to $O(h^{2})$ for all $p\in(2,\infty)$ and $q\in[2,\infty)$.
  3. 3.
    The authors in [28] derived a strong convergence rate for a spatial semi-discretization of Eq. 1.1 with a rough initial value $v\in L^{\infty}$. However, to the best of our knowledge, there is currently no strong convergence estimate for fully discrete approximations of the three-dimensional stochastic Allen-Cahn equation with a rough initial value. It would be of interest to establish similar error estimates as Eq. 3.2 for fully discrete approximations with rough initial values.
4 Proofs
In this section, we adopt the following notational conventions:
[…]
  • For $\alpha\in\mathbb{R}$ and $q\in(1,\infty)$, $\dot{H}_{h}^{\alpha,q}$ represents the Banach space $X_{h}$ equipped with the norm
    $\lVert{u_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}}:=\lVert{(-\Delta_{h})^{\alpha/2}u_{h}}\rVert_{L^{q}},\quad\forall u_{h}\in X_{h}.$
  • The symbol $c$ denotes a generic positive constant, which is independent of the spatial mesh size $h$ and the time step $\tau$. Its specific value may depend on the regularity parameters associated with the triangulation $\mathcal{K}_{h}$, and may change from line to line.
The subsequent structure of this section is outlined as follows. In Subsection 4.1, we introduce several preliminary lemmas that form the foundation for our subsequent analysis. Subsequently, Subsection 4.2 is dedicated to establishing a stability estimate for a discrete stochastic convolution, which is of independent interest and also crucial for the proof of Theorem 3.1. Lastly, Subsection 4.3 presents a detailed proof of Theorem 3.1.
4.1 Preliminary lemmas
According to the theoretical results in [9], the projection operator $P_{h}$ has the following well-known stability property.
Lemma 4.1.
For any $\alpha\in[0,2]$ and $q\in(1,\infty)$, $\lVert{P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},\,\dot{H}_{h}^{\alpha,q})}$ is uniformly bounded with respect to the spatial mesh size $h$.
Referring to the established estimates for the resolvents of $\Delta_{h}$ as established in [2], a standard computation (for which the relevant techniques can be found in [44, Section 7.7 in Chapter 2]) yields the following well-recognized inequalities.
Lemma 4.2.
Let $q\in(1,\infty)$ and $0\leqslant\alpha\leqslant\beta<\infty$. Then the following inequalities hold:
$\lVert{e^{t\Delta_{h}}}\rVert_{\mathcal{L}(\dot{H}_{h}^{\alpha,q},\dot{H}_{h}^{\beta,q})}\leqslant ct^{(\alpha-\beta)/2},\quad\forall t\in(0,T];$ (4.1)
$\lVert{I-e^{t\Delta_{h}}}\rVert_{\mathcal{L}(\dot{H}_{h}^{\beta,q},\dot{H}_{h}^{\alpha,q})}\leqslant ct^{(\beta-\alpha)/2},\quad\forall t\in[0,T];$ (4.2)
$\lVert{(I-\tau\Delta_{h})^{-1}-I}\rVert_{\mathcal{L}(\dot{H}_{h}^{\beta,q},\dot{H}_{h}^{\alpha,q})}\leqslant c\tau^{(\beta-\alpha)/2};$ (4.3)
$\lVert{(I-\tau\Delta_{h})^{-m}}\rVert_{\mathcal{L}(\dot{H}_{h}^{\alpha,q},\dot{H}_{h}^{\beta,q})}\leqslant\frac{c}{(m\tau)^{(\beta-\alpha)/2}},\quad\,\forall m\in\mathbb{N}_{>0}.$ (4.4)
We have the following fundamental estimate for the stochastic integrals; see, e.g., [42, pp. 792].
Lemma 4.3.
Let $p\in(1,\infty)$ and $q\in[2,\infty)$. There exists a positive constant $c$ such that for any $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$, the following inequality holds:
$\Big{\lVert}{\int_{0}^{T}g(t)\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{p}(\Omega;L^{q})}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega;L^{2}(0,T;\gamma(H;L^{q})))}.$
Lemma 4.4.
Suppose $q\in[2,\infty)$. The following inequalities hold:
$\lVert{F(u)}\rVert_{\gamma(H,L^{q})}\leqslant c(1+\lVert{u}\rVert_{L^{q}})$ for all $\text{for all }u\in L^{q},$ (4.5)
$\lVert{F(u)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c(1+\lVert{u}\rVert_{\dot{H}^{1,q}})$ for all $\text{for all }u\in\dot{H}^{1,q},$ (4.6)
$\lVert{F(u_{1})-F(u_{2})}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u_{1}-u_{2}}\rVert_{L^{q}}$ for all $\text{for all }u_{1},u_{2}\in L^{q}.$ (4.7)
Proof.
These properties are derived from straightforward calculations based on the definition of $F$. For a detailed derivation, see Lemma 4.5 in [28]. ∎
We have the following discrete deterministic maximal $L^{p}$-regularity estimate; see [20, 26].
Lemma 4.5.
Let $p,q\in(1,\infty)$. Assume that $(Z_{j})_{j=0}^{J}$ is the solution of the discretization
$Z_{j+1}-Z_{j}=\tau\Delta_{h}Z_{j+1}+P_{h}\int_{t_{j}}^{t_{j+1}}g(t)\,\mathrm{d}t,\quad 0\leqslant j<J,$
$Z_{0}=0,$
where $g\in L^{p}(0,T;L^{q})$. Then the following stability estimate holds:
$\Big{(}\tau\sum_{j=1}^{J}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{2,q}}^{p}\Big{)}^{1/p}\leqslant c\lVert{g}\rVert_{L^{p}(0,T;L^{q})}.$
Finally, we present some standard properties of the space $\dot{H}_{h}^{\alpha,q}$, where $\alpha\in\mathbb{R}$ and $q\in(1,\infty)$. For any $q\in(1,\infty)$, there exist two positive constants $c_{0}$ and $c_{1}$, independent of $h$, such that for all $u_{h}\in X_{h}$, the following inequality holds:
$c_{0}\lVert{u_{h}}\rVert_{\dot{H}^{1,q}}\leqslant\lVert{u_{h}}\rVert_{\dot{H}_{h}^{1,q}}\leqslant c_{1}\lVert{u_{h}}\rVert_{\dot{H}^{1,q}}.$
For any $\alpha\in(0,2)$ and $q\in(1,\infty)$, there exist two positive constants $c_{2}$ and $c_{3}$, independent of $h$, such that for all $u_{h}\in X_{h}$, the following inequality holds:
$c_{2}\lVert{u_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}}\leqslant\lVert{u_{h}}\rVert_{[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}}\leqslant c_{3}\lVert{u_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}},$
where $[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}$ is the complex interpolation space between $\dot{H}_{h}^{0,q}$ and $\dot{H}_{h}^{2,q}$ (see [44, Section 5 in Chapter 1]). Furthermore, for any $\beta\in[3/2,\infty)$ and $q\in[2,\infty)$, using the well-known estimate (see, e.g., [6, Section 8.5]),
$\lVert{\Delta^{-1}-\Delta_{h}^{-1}P_{h}}\rVert_{\mathcal{L}(L^{\beta})}\leqslant ch^{2},$
combined with the continuous embedding of $\dot{H}^{2,\beta}$ into $L^{q}$ and the inverse estimate, it follows that the continuous embedding of $\dot{H}_{h}^{2,\beta}$ into $L^{q}$ is independent of $h$. These results will be used throughout the subsequent analysis without further explicit reference.
4.2 A stability estimate for a discrete stochastic convolution
The objective of this subsection is to establish the stability estimate
$\bigg{(}\mathbb{E}\Big{[}\max_{1\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}\bigg{)}^{1/p}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))}$ (4.8)
for the discrete stochastic convolution defined by
$Z_{j}=\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}(I-\tau\Delta_{h})^{k-j}g_{h}(t)\,\mathrm{d}W_{H}(t),\quad 1\leqslant j\leqslant J,$ (4.9)
where $g_{h}$ belongs to $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))$, with parameters $p\in(2,\infty)$ and $q\in[2,\infty)$. This stability estimate is crucial for the derivation of pathwise uniform convergence rates for the numerical approximations of a broad class of nonlinear stochastic parabolic equations.
Within the framework of Hilbert spaces, we note that Gyöngy and Millet [13, Theorem 2.6] have established the following stability estimate:
$\left(\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\|Z_{j}\|_{L^{2}}^{2}\right]\right)^{1/2}\leqslant c\left\|g\right\|_{L^{2}(\Omega\times(0,T);\gamma(H,L^{2}))}.$
Moreover, under the Banach space setting, a significant advancement was made by van Neerven and Veraar [40, Proposition 5.4], who demonstrated the following sharp estimate for a discrete stochastic convolution without spatial discretization:
$\left(\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\Big{\|}\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}(I-\tau\Delta)^{k-j}g(t)\,\mathrm{d}W_{H}(t)\Big{\|}_{L^{q}}^{p}\right]\right)^{1/p}\leqslant c\left\|g\right\|_{L^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))},$
holds for all $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$ with $p,q\in[2,\infty)$. The derivation of this estimate relies on the contraction property of the operator $(I-\tau\Delta)^{-1}$ on $L^{q}$ for all $q\in[2,\infty)$. However, it is important to note that $(I-\tau\Delta_{h})^{-1}$ fails to possess the contraction property on $\dot{H}_{h}^{0,q}$ for any $q\in(2,\infty)$. Hence, Proposition 5.4 from [42] can not be directly applied to bound the discrete stochastic convolution defined by Eq. 4.9.
To address the aforementioned challenge, we employ the discrete stochastic maximal $L^{p}$-regularity estimate for
$\tau\sum_{j=1}^{J}\mathbb{E}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p},\quad p\in(2,\infty),\,q\in[2,\infty),$
coupled with the approximation property of $(I-\tau\Delta_{h})^{-1}$ to the identity operator $I$ as presented in Eq. 4.3. We formally state the discrete stochastic maximal $L^{p}$-regularity estimate in the following proposition.
Proposition 4.1.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Assume that $(Z_{j})_{j=1}^{J}$ is defined by Eq. 4.9, where $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))$. Then the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
$\Bigg{(}\mathbb{E}\bigg{[}\tau\sum_{j=1}^{J}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}\Bigg{)}^{1/p}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))}.$
Proof.
Note that [29, Theorem 3.1] has demonstrated that the $H^{\infty}$-calculus of $-\Delta_{h}$ is uniformly bounded with respect to the spatial mesh size $h$. Subsequently, applying [27, Theorem 3.2], we establish the desired discrete stochastic maximal $L^{p}$-regularity estimate. ∎
Now we present the main result of this subsection in the form of the following proposition.
Proposition 4.2.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Suppose that $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))$, and let $(Z_{j})_{j=1}^{J}$ be defined by Eq. 4.9. Then the stability estimate Eq. 4.8 holds.
Proof.
Let $Z_{0}:=0$ and, for any $t\in[0,T]$, define
$\Psi(t):=(I-\tau\Delta_{h})^{-1}g_{h}(t),\quad\mathcal{G}(t):=p\sum_{j=0}^{J-1}\mathbbm{1}_{[t_{j},t_{j+1})}(t)\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\left\langle\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\Psi(t)\right\rangle,$
where $\mathbbm{1}_{[t_{j},t_{j+1})}$ is the indicator function for the time interval $[t_{j},t_{j+1})$. For any $0\leqslant j<J$, we also define
$\delta_{j}:=(I-\tau\Delta_{h})^{-1}Z_{j}-Z_{j},\quad M_{j}:=\delta_{j}+\int_{t_{j}}^{t_{j+1}}\Psi(t)\,\mathrm{d}W_{H}(t).$
It can be verified that
$Z_{j+1}=Z_{j}+M_{j}\quad\text{for all $0\leqslant j<J$}.$
The rest of the proof is divided into the following five steps.
Step 1. Let us prove
$\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}\leqslant I_{1}+I_{2}+I_{3}+I_{4},$ (4.10)
where
$\hskip-165.00024ptI_{1}:=p\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\left|\left\langle\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\delta_{j}\right\rangle\right|\bigg{]},$
$\hskip-195.00029ptI_{2}:=\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\Big{\lvert}{\int_{0}^{t}\mathcal{G}(s)\,\mathrm{d}W_{H}(s)}\Big{\rvert}\bigg{]},$
$I_{3}:=p|p-q|\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\big{\rangle}^{2}\,\mathrm{d}\theta\bigg{]},$
$\hskip-60.00009ptI_{4}:=p(q-1)\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2},\,M_{j}^{2}}\big{\rangle}\,\mathrm{d}\theta\bigg{]}.$
Fix any $0\leqslant j<J$. Define the function $g:[0,1]\to\mathbb{R}$ by
$g(\theta):=\big{\lVert}{Z_{j}+\theta M_{j}}\big{\rVert}_{L^{q}}^{p},\quad\theta\in[0,1].$
A straightforward computation yields
$g^{\prime}(0)=p\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,M_{j}}\big{\rangle}$
$\qquad=p\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\delta_{j}}\big{\rangle}+\int_{t_{j}}^{t_{j+1}}\mathcal{G}(t)\,\mathrm{d}W_{H}(t)$
and
$g^{\prime\prime}(\theta)=p(p-q)\big{\lVert}{Z_{j}+\theta M_{j}}\big{\rVert}_{L^{q}}^{p-2q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\big{\rangle}^{2}+{}$
$\qquad\qquad\quad p(q-1)\big{\lVert}{Z_{j}+\theta M_{j}}\big{\rVert}_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2},\,M_{j}^{2}}\big{\rangle},\quad\forall\theta\in[0,1].$
Hence, by $g(1)=g(0)+g^{\prime}(0)+\int_{0}^{1}(1-\theta)g^{\prime\prime}(\theta)\mathrm{d}\theta$ and the fact that $g(1)=\lVert{Z_{j+1}}\rVert_{L^{q}}^{p}$ and $g(0)=\lVert{Z_{j}}\rVert_{L^{q}}^{p}$, we obtain
$\lVert{Z_{j+1}}\rVert_{L^{q}}^{p}$ $\leqslant\lVert{Z_{j}}\rVert_{L^{q}}^{p}+p\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\delta_{j}}\big{\rangle}+\int_{t_{j}}^{t_{j+1}}\mathcal{G}(t)\,\mathrm{d}W_{H}(t)$
$\quad{}+p|p-q|\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\big{\rangle}^{2}\,\mathrm{d}\theta$
$\quad{}+p(q-1)\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}+\theta M_{j}}\rvert^{q-2},\,M_{j}^{2}}\big{\rangle}\,\mathrm{d}\theta.$
Given that $Z_{0}=0$ and the preceding inequality holds for all $0\leqslant j<J$, we deduce the following bound for the supremum of the sequence $(Z_{j})_{j=0}^{J}$:
$\max_{0\leqslant j\leqslant J}\left\|Z_{j}\right\|_{L^{q}}^{p}\leqslant p\sum_{j=0}^{J-1}\left\|Z_{j}\right\|_{L^{q}}^{p-q}\left|\big{\langle}{\left|Z_{j}\right|^{q-2}Z_{j},\,\delta_{j}}\big{\rangle}\right|+\sup_{t\in[0,T]}\left|\int_{0}^{t}\mathcal{G}(s)\,\mathrm{d}W_{H}(s)\right|$
$\qquad\qquad{}+p|p-q|\sum_{j=0}^{J-1}\int_{0}^{1}\left\|Z_{j}+\theta M_{j}\right\|_{L^{q}}^{p-2q}\big{\langle}{\left|Z_{j}+\theta M_{j}\right|^{q-2}(Z_{j}+\theta M_{j}),\,M_{j}}\big{\rangle}^{2}\,\mathrm{d}\theta$
$\qquad\qquad{}+p(q-1)\sum_{j=0}^{J-1}\int_{0}^{1}\left\|Z_{j}+\theta M_{j}\right\|_{L^{q}}^{p-q}\big{\langle}{\left|Z_{j}+\theta M_{j}\right|^{q-2},\,M_{j}^{2}}\big{\rangle}\,\mathrm{d}\theta.$
By taking expectations on both sides of the inequality, we arrive at the desired inequality (4.10).
Step 2. From the definition of $I_{1}$, it follows that
$I_{1}$ $\leqslant p\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\lVert}{\lvert{Z_{j}}\rvert^{q-2}Z_{j}}\big{\rVert}_{\dot{H}^{1,q/(q-1)}}\lVert{\delta_{j}}\rVert_{\dot{H}^{-1,q}}\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\lVert}{\lvert{Z_{j}}\rvert^{q-2}Z_{j}}\big{\rVert}_{\dot{H}^{1,q/(q-1)}}\lVert{\delta_{j}}\rVert_{\dot{H}_{h}^{-1,q}}\bigg{]}.$
For any $0\leqslant j<J$, we observe that
$\|\delta_{j}\|_{\dot{H}_{h}^{-1,q}}$ $\leqslant\|(I-\tau\Delta_{h})^{-1}-I\|_{\mathcal{L}(\dot{H}_{h}^{1,q},\dot{H}_{h}^{-1,q})}\|Z_{j}\|_{\dot{H}_{h}^{1,q}}$
$=\|(I-\tau\Delta_{h})^{-1}-I\|_{\mathcal{L}(\dot{H}_{h}^{2,q},\dot{H}_{h}^{0,q})}\|Z_{j}\|_{\dot{H}_{h}^{1,q}}$
$\leqslant c\tau\|Z_{j}\|_{\dot{H}_{h}^{1,q}},$
where the last inequality is justified by (4.3) with $\alpha=0$ and $\beta=2$. Furthermore,
$\|\,|Z_{j}|^{q-2}Z_{j}\|_{\dot{H}^{1,q/(q-1)}}$ $\leqslant c\|\nabla(|Z_{j}|^{q-2}Z_{j})\|_{L^{q/(q-1)}}$
[…]
$\leqslant c\|Z_{j}\|_{\dot{H}_{h}^{1,q}}^{q-1}.$
Consequently, putting the above estimates together, we obtain
$I_{1}\leqslant c\mathbb{E}\bigg{[}\tau\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}.$ (4.11)
Step 3. By the Burkholder-Davis-Gundy inequality and Hölder’s inequality, we obtain
$I_{2}$ $\leqslant c\mathbb{E}\Big{[}\int_{0}^{T}\lVert{\mathcal{G}(t)}\rVert_{\gamma(H,\mathbb{R})}^{2}\,\mathrm{d}t\Big{]}^{1/2}$
$=c\mathbb{E}\Big{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\Big{\lVert}{\lVert{Z_{j}}\rVert_{L^{q}}^{p-q}\big{\langle}{\lvert{Z_{j}}\rvert^{q-2}Z_{j},\,\Psi(t)}\big{\rangle}}\Big{\rVert}_{\gamma(H,\mathbb{R})}^{2}\,\mathrm{d}t\Big{]}^{1/2}$
$\leqslant c\mathbb{E}\Big{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{Z_{j}}\rVert_{L^{q}}^{2p-2q}\big{\lVert}{\lvert{Z_{j}}\rvert^{q-2}Z_{j}}\big{\rVert}_{L^{q/(q-1)}}^{2}\lVert{\Psi(t)}\rVert_{\gamma(H,L^{q})}^{2}\,\mathrm{d}t\Big{]}^{1/2}$
$=c\mathbb{E}\Big{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{Z_{j}}\rVert_{L^{q}}^{2p-2}\lVert{\Psi(t)}\rVert_{\gamma(H,L^{q})}^{2}\,\mathrm{d}t\Big{]}^{1/2}.$
It follows that
$I_{2}\leqslant c\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p-1}\lVert{\Psi}\rVert_{L^{2}(0,T;\gamma(H,L^{q}))}\Big{]}.$
By Young’s inequality, we then obtain
$I_{2}\leqslant\frac{1}{2}\mathbb{E}\Big{[}\max_{0\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\Big{]}+c\lVert{\Psi}\rVert_{L^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))}^{p}.$ (4.12)
Step 4. We employ Hölder’s inequality and Young’s inequality as follows:
$I_{3}+I_{4}$ $\leqslant c\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\int_{0}^{1}\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p-2}\lVert{M_{j}}\rVert_{L^{q}}^{2}\,\mathrm{d}\theta\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\mathbb{E}\Big{[}\int_{0}^{1}\tau\lVert{Z_{j}+\theta M_{j}}\rVert_{L^{q}}^{p}+\tau^{1-p/2}\lVert{M_{j}}\rVert_{L^{q}}^{p}\,\mathrm{d}\theta\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\tau\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\bigg{]}+c\tau^{1-p/2}\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{M_{j}}\rVert_{L^{q}}^{p}\bigg{]}.$
Moreover,
$\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{M_{j}}\rVert_{L^{q}}^{p}\bigg{]}$ $\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{\delta_{j}}\rVert_{L^{q}}^{p}\bigg{]}+c\tau^{p/2-1}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{\Psi(t)}\rVert_{L^{p}(\Omega;\gamma(H,L^{q}))}^{p}\,\mathrm{d}t$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\tau^{p/2}\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}+c\tau^{p/2-1}\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p},$
where in the first inequality (i), we invoke the definition $M_{j}:=\delta_{j}+\int_{t_{j}}^{t_{j+1}}\Psi(t)\,\mathrm{d}W_{H}(t)$, Lemma 4.3, and Hölder’s inequality; and in the second inequality (ii), we make use of the definition $\delta_{j}:=(I-\tau\Delta_{h})^{-1}Z_{j}-Z_{j}$, along with the resolvent estimate from Lemma 4.3 with $\alpha=0$ and $\beta=1$. By combining these estimates, we obtain
$I_{3}+I_{4}\leqslant c\mathbb{E}\bigg{[}\tau\sum_{j=0}^{J-1}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}+c\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$ (4.13)
Step 5. Combining Eqs. 4.10, 4.11, 4.12 and 4.13, and noting that $Z_{0}=0$, we deduce that
$\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\lVert{Z_{j}}\rVert_{L^{q}}^{p}\right]\leqslant c\mathbb{E}\bigg{[}\tau\sum_{j=1}^{J-1}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}+c\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}^{p}.$
Applying Proposition 4.1, we infer that
$\mathbb{E}\bigg{[}\tau\sum_{j=1}^{J-1}\lVert{Z_{j}}\rVert_{\dot{H}_{h}^{1,q}}^{p}\bigg{]}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))}^{p}.$
Moreover, using inequality (4.4) with parameters $\alpha=\beta=0$, we establish
$\lVert{\Psi}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))}.$
Aggregating these estimates, we conclude the desired stability estimate (4.8). This completes the proof. ∎
Remark 4.1.
Considering the maximal inequality established in [40, Theorem 1.1],
$\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\left\lVert\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s)\right\rVert_{L^{q}}^{p}\bigg{]}\leqslant c\mathbb{E}\left(\int_{0}^{T}\lVert g(t)\rVert_{\gamma(H,L^{q})}^{2}\,\mathrm{d}t\right)^{p/2},$
for all $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$ with $p,q\in[2,\infty)$, the stability estimate provided in Proposition 4.2 is suboptimal. While Proposition 4.2 is adequate for the purposes of the current investigation, it is an interesting question to derive a sharper stability estimate than the one given by (4.8).
4.3 Proof of Theorem 3.1
In this subsection, we always assume that the conditions of Theorem 3.1 are satisfied. Let $y_{h}$ be the strong solution of the following spatial semi-discretization:
$\begin{cases}\mathrm{d}y_{h}(t)=(\Delta_{h}y_{h}+y_{h}-P_{h}y_{h}^{3})(t)\,\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),&t\in[0,T],\\ y_{h}(0)=P_{h}v.\end{cases}$ (4.14)
Li and Zhou [28, Theorem 3.3] have established the error estimate
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\varepsilon}$ (4.15)
for all $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$. Therefore, to complete the proof of Theorem 3.1, it remains to prove
$\Big{\lVert}{\max_{1\leqslant j\leqslant J}\lVert{y_{h}(t_{j})-Y_{j}}\rVert_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}\leqslant c\tau^{1/2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty),$ (4.16)
which is the primary focus of the remainder of this subsection.
To this end, we first note that by the regularity property of the mild solution $y$ as stated in Proposition 2.1, the stability property of $P_{h}$ as stated in Lemma 4.1, the standard approximation properties of $P_{h}$, the error estimate (4.15), the growth estimate of $F$ in (4.6), and the inverse estimate, we deduce the following stability assertion for $y_{h}$:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{2-\epsilon,q}))}+\lVert{P_{h}y_{h}^{3}}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))}+\lVert{P_{h}F(y_{h})}\rVert_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{1,q}))}$ (4.17)
[…]
With this stability assertion and the inequalities (4.1) and (4.2), a routine calculation yields the following estimates:
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j})}\rVert_{L^{q}}^{p}\,\mathrm{d}t\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty),$ (4.18)
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{L^{q}}^{p}\,\mathrm{d}t\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty),$ (4.19)
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1/2+1/p,q}}^{p}\,\mathrm{d}t\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty),$ (4.20)
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\rVert_{L^{q}}^{p}\,\mathrm{d}t\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.21)
The structure of the remaining proof is outlined as follows:
  1. 1.
    In Subsection 4.3.1, we derive the error estimate
    $\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\lVert{y_{h}(t_{j})-Y_{j}}\rVert_{L^{2}}^{p}\biggr{]}\leqslant c\tau^{p/2},\quad\forall p\in[2,\infty).$ (4.22)
  2. 2.
    In Subsection 4.3.2, we establish the following stability property:
    $\mathbb{E}\bigg{[}\tau\sum_{j=1}^{J}\lVert{Y_{j}}\rVert_{L^{q}}^{p}\biggr{]}\text{ is uniformly bounded w.r.t.~{}$h$ and $\tau$ for all $p,q\in(2,\infty)$}.$ (4.23)
  3. 3.
    In Subsection 4.3.3, we establish the error estimate
    $\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\lVert{Y_{j}-y_{h}(t_{j})}\rVert_{L^{q}}^{p}\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.24)
  4. 4.
    Finally, in Subsection 4.3.4, by utilizing Eqs. 4.17, 4.18, 4.19, 4.20, 4.21, 4.22, 4.23 and 4.24, along with the discrete deterministic maximal $L^{p}$-regularity estimate from Lemma 4.5, the discrete stochastic maximal $L^{p}$-regularity estimate from Proposition 4.1, and the stability estimate from Proposition 4.2, we conclude the proof for the desired error estimate (4.16).
4.3.1 Proof of the error estimate Eq. 4.22
For each $0\leqslant j\leqslant J$, let $e_{j}:=Y_{j}-y_{h}(t_{j})$. It is clear that
$e_{j+1}-e_{j}$ $=\int_{t_{j}}^{t_{j+1}}\Delta_{h}\big{[}Y_{j+1}-y_{h}(t)\big{]}+Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\,\mathrm{d}t$ (4.25)
$\quad{}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad\forall 0\leqslant j<J.$
We introduce the decomposition $e_{j}=\xi_{j}+\eta_{j}$ for each $0\leqslant j\leqslant J$, defining the sequence $(\xi_{j})_{j=0}^{J}$ as follows:
$\begin{cases}\xi_{j+1}-\xi_{j}=\tau\Delta_{h}\xi_{j+1}+\int_{t_{j}}^{t_{j+1}}\Delta_{h}\big{[}y_{h}(t_{j+1})-y_{h}(t)\big{]}\,\mathrm{d}t,\quad 0\leqslant j<J,\\ \xi_{0}=0.\end{cases}$
It can be readily verified that
$\eta_{j+1}-\eta_{j}$ $=\int_{t_{j}}^{t_{j+1}}\left[\Delta_{h}\eta_{j+1}+Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\right]\,\mathrm{d}t$ (4.26)
$\quad{}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad\forall 0\leqslant j<J.$
Moreover, by Lemma 4.5 and Eq. 4.19 we obtain
$\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\lVert{\xi_{j}}\rVert_{L^{q}}^{p}\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.27)
The remainder of the proof is organized into the following four steps.
Step 1. Let us prove that, for any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$,
$\mathbb{E}\lVert{\eta_{j+1}}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\mathbb{E}\Big{\lVert}{\eta_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{2}}^{2m}$ (4.28)
$\quad{}+c\tau\mathbb{E}\lVert{\eta_{j}}\rVert_{L^{2}}^{2m}+c\tau\mathbb{E}\lVert{\xi_{j}}\rVert_{L^{2}}^{2m}+c\mathbb{E}\big{[}A_{j,m}+B_{j,m}\big{]},$
where
$A_{j,m}$ $:=\tau\langle{\xi_{j+1}^{2},\,y_{h}^{2}(t_{j+1})+Y_{j+1}^{2}}\rangle^{m},$
$B_{j,m}$ $:=\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j})}\rVert_{L^{2}}^{2m}+\lVert{y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\rVert_{L^{2}}^{2m}\,\mathrm{d}t.$
To this end, fix any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$. Consider the function
$g(\theta)$ $:=\eta_{j}+\theta\int_{t_{j}}^{t_{j+1}}\Big{[}\Delta_{h}\eta_{j+1}+Y_{j}-y_{h}(t)+P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\Big{]}\,\mathrm{d}t$
$\qquad{}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad\theta\in[0,1].$
For any $\theta\in[0,1]$, differentiating $\lVert{g(\theta)}\rVert_{L^{2}}^{2m}$ with respect to $\theta$ yields
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}$ $=2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\langle{g(\theta),g^{\prime}(\theta)}\rangle$
$=2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\langle{g(1)-(1-\theta)g^{\prime}(\theta),g^{\prime}(\theta)}\rangle$
$\leqslant 2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\langle{g(1),g^{\prime}(\theta)}\rangle.$
Since $g(1)=\eta_{j+1}$, we have for any $\theta\in[0,1]$ that
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}\leqslant 2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\biggl{[}\int_{t_{j}}^{t_{j+1}}\langle{\eta_{j+1},\,y_{h}^{3}(t_{j+1})-Y_{j+1}^{3}}\rangle\,\mathrm{d}t$
$+\int_{t_{j}}^{t_{j+1}}\big{\langle}{\eta_{j+1},\,\Delta_{h}\eta_{j+1}+Y_{j}-y_{h}(t)+y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\big{\rangle}\,\mathrm{d}t\biggr{]}.$
Using the fact $Y_{i}-y_{h}(t_{i})=\xi_{i}+\eta_{i}$ for all $0\leqslant i\leqslant J$, an elementary calculation yields
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}\leqslant 2m\lVert{g(\theta)}\rVert_{L^{2}}^{2m-2}\biggl{[}\int_{t_{j}}^{t_{j+1}}\langle{\xi_{j+1}^{2},\,y_{h}^{2}(t_{j+1})+Y_{j+1}^{2}}\rangle\,\mathrm{d}t$
$\qquad\quad{}+\int_{t_{j}}^{t_{j+1}}\lVert{\eta_{j}}\rVert_{L^{2}}^{2}+\lVert{\xi_{j}}\rVert_{L^{2}}^{2}+\lVert{y_{h}(t)-y_{h}(t_{j})}\rVert_{L^{2}}^{2}+\lVert{y_{h}^{3}(t)-y_{h}^{3}(t_{j+1})}\rVert_{L^{2}}^{2}\,\mathrm{d}t\biggr{]}.$
Hence, applying Young’s inequality and using the definitions of $A_{j,m}$, and $B_{j,m}$, we obtain the following differential inequality: for any $\theta\in[0,1]$,
$\frac{\mathrm{d}}{\mathrm{d}\theta}\lVert{g(\theta)}\rVert_{L^{2}}^{2m}$ $\leqslant c\tau\lVert{g(\theta)}\rVert_{L^{2}}^{2m}+c\tau\lVert{\eta_{j}}\rVert_{L^{2}}^{2m}+c\tau\lVert{\xi_{j}}\rVert_{L^{2}}^{2m}+c\big{(}A_{j,m}+B_{j,m}\big{)}.$
Employing Gronwall’s lemma subsequently gives
$\lVert{g(1)}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\lVert{g(0)}\rVert_{L^{2}}^{2m}+c\tau\lVert{\eta_{j}}\rVert_{L^{2}}^{2m}+c\tau\lVert{\xi}\rVert_{L^{2}}^{2m}+c\left(A_{j,m}+B_{j,m}\right).$
Taking the expectations on both sides yields
$\mathbb{E}\lVert{g(1)}\rVert_{L^{2}}^{2m}\leqslant(1+c\tau)\mathbb{E}\lVert{g(0)}\rVert_{L^{2}}^{2m}+c\tau\mathbb{E}\lVert{\eta_{j}}\rVert_{L^{2}}^{2m}+c\tau\mathbb{E}\lVert{\xi_{j}}\rVert_{L^{2}}^{2m}+c\mathbb{E}\big{[}A_{j,m}+B_{j,m}\big{]}.$
The desired claim Eq. 4.28 then follows by recognizing that $g(1)=\eta_{j+1}$ and $g(0)=\eta_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)$.
Step 2. Fix any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$. Using Hölder’s inequality, Young’s inequality, and the fact
$\mathbb{E}\biggl{[}\lVert{\eta_{j}}\rVert_{L^{2}}^{2m-2}\Big{\langle}{\eta_{j},\,\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rangle}\biggr{]}=0,$
a straightforward calculation yields
$\mathbb{E}\Big{\lVert}{\eta_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{2}}^{2m}$
[…]
For the second term on the right-hand side of this inequality, we can use the fact that $P_{h}$ is the $L^{2}$-orthogonal projection operator onto $X_{h}$. Additionally, we apply Lemma 4.3, Hölder’s inequality, and the Lipschitz continuity of $F$ from Eq. 4.7 to deduce that
$\tau^{1-m}\mathbb{E}\Big{\lVert}{P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{2}}^{2m}\leqslant c\int_{t_{j}}^{t_{j+1}}\mathbb{E}\lVert{Y_{j}-y_{h}(t)}\rVert_{L^{2}}^{2m}\,\mathrm{d}t.$
Consequently,
$\mathbb{E}\Big{\lVert}{\eta_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{2}}^{2m}$
[…]
Using the fact $Y_{j}-y_{h}(t_{j})=\xi_{j}+\eta_{j}$, along with the definition of $B_{j,m}$, we readily conclude that, for any $m\in\mathbb{N}_{>0}$ and $0\leqslant j<J$,
$\mathbb{E}\Big{\lVert}{\eta_{j}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t)}\Big{\rVert}_{L^{2}}^{2m}$ (4.29)
[…]
Step 3. By combining Eq. 4.28 and Eq. 4.29, we arrive at the conclusion that for any $m\in\mathbb{N}_{>0}$ and for all $0\leqslant j<J$,
$\|\eta_{j+1}\|_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant(1+c\tau)\|\eta_{j}\|_{L^{2m}(\Omega;L^{2})}^{2m}+c\tau\mathbb{E}\lVert{\xi_{j}}\rVert_{L^{2}}^{2m}+c\mathbb{E}\big{[}A_{j,m}+B_{j,m}\big{]}.$
Applying the discrete Gronwall’s lemma and recalling that $\eta_{0}=0$, we deduce that
$\max_{0\leqslant j\leqslant J}\|\eta_{j}\|_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant c\sum_{j=0}^{J-1}\mathbb{E}(\tau\lVert{\xi_{j}}\rVert_{L^{2}}^{2m}+A_{j,m}+B_{j,m}),\quad\forall m\in\mathbb{N}_{>0}.$
Taking into account the inequalities Eqs. 4.18, 4.21 and 4.27, we then obtain that
$\max_{0\leqslant j\leqslant J}\|\eta_{j}\|_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant c\tau^{m}+c\mathbb{E}\Big{[}\sum_{j=0}^{J-1}A_{j,m}\Big{]},\quad\forall m\in\mathbb{N}_{>0}.$ (4.30)
Step 4. Employing an analogous argument to that in Steps 1 and 2, and using the stability property of $P_{h}$ as stated in Lemma 4.1, as well as the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we can conclude that (see, also, [33, Lemma 4.1]) for any $m\in\mathbb{N}_{>0}$,
$\mathbb{E}\Bigl{[}\tau\sum_{j=0}^{J}\lVert{Y_{j}}\rVert_{L^{2}}^{2m}\Bigr{]}+\mathbb{E}\Bigl{[}\tau\sum_{j=0}^{J}\lVert{Y_{j}}\rVert_{L^{4}}^{4}\Bigr{]}$
is uniformly bounded with respect to $h$ and $\tau$. By Hölder’s inequality, we further infer that for any $p_{0}\in[2,\infty)$, there exists $q_{0}\in(2,\infty)$ such that
$\mathbb{E}\Bigl{[}\tau\sum_{j=1}^{J}\lVert{Y_{j}}\rVert_{L^{q_{0}}}^{p_{0}}\Bigr{]}$
is uniformly bounded with respect to $h$ and $\tau$. In view of this assertion and Eq. 4.27, we can use Hölder’s inequality to infer that
$\mathbb{E}\Big{[}\tau\sum_{j=0}^{J-1}\langle{\xi_{j+1}^{2},\,Y_{j+1}^{2}}\rangle^{m}\Big{]}\leqslant c\tau^{m},\quad\forall m\in\mathbb{N}_{>0}.$
Similarly, using the stability estimate of $y_{h}$ in Eq. 4.17, we have
$\mathbb{E}\Big{[}\tau\sum_{j=0}^{J-1}\langle{\xi_{j+1}^{2},\,y_{h}^{2}(t_{j+1})}\rangle^{m}\Big{]}\leqslant c\tau^{m},\quad\forall m\in\mathbb{N}_{>0}.$
Consequently,
$\mathbb{E}\Big{[}\sum_{j=0}^{J-1}A_{j,m}\Big{]}\leqslant c\tau^{m},\quad\forall m\in\mathbb{N}_{>0},$
which, combined with Eq. 4.30, further implies that
$\max_{0\leqslant j\leqslant J}\lVert{\eta_{j}}\rVert_{L^{2m}(\Omega;L^{2})}^{2m}\leqslant c\tau^{m},\quad\forall m\in\mathbb{N}_{>0}.$
Combining this inequality with Eq. 4.27, and noting that $Y_{j}-y_{h}(t_{j})=\xi_{j}+\eta_{j}$ for all $0\leqslant j\leqslant J$, we confirm the validity of Eq. 4.22 for all $p=2m$, where $m$ is a positive integer. The generalization to $p\in[2,\infty)$ is achieved by invoking Hölder’s inequality. This completes the proof of the error estimate Eq. 4.22.
4.3.2 Proof of the stability property Eq. 4.23
From Eq. 3.1, we decompose the numerical solution into three components:
$Y_{j}=\eta_{j}^{(0)}+\eta_{j}^{(1)}+\eta_{j}^{(2)}\quad\text{for all }0\leqslant j\leqslant J.$
These components are defined by the recursive relations:
$\left\{\begin{array}[]{ll}\eta^{(0)}_{j+1}-\eta^{(0)}_{j}=\tau\Delta_{h}\eta^{(0)}_{j+1},&\quad 0\leqslant j<J,\\ \eta_{0}^{(0)}=P_{h}v,\end{array}\right.$
$\left\{\begin{array}[]{ll}\eta^{(1)}_{j+1}-\eta^{(1)}_{j}=\tau\Delta_{h}\left(\eta^{(1)}_{j+1}+Y_{j}-P_{h}Y_{j+1}^{3}\right),&\quad 0\leqslant j<J,\\ \eta_{0}^{(1)}=0,\end{array}\right.$
and
$\left\{\begin{array}[]{ll}\eta^{(2)}_{j+1}-\eta^{(2)}_{j}=\tau\Delta_{h}\eta^{(2)}_{j+1}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})\,\mathrm{d}W_{H}(t),&\quad 0\leqslant j<J,\\ \eta_{0}^{(2)}=0.\end{array}\right.$
Given the identity $\eta_{j}^{(0)}=(I-\tau\Delta_{h})^{-j}P_{h}v$ for $0\leqslant j\leqslant J$, and using the stability property of $P_{h}$ from Lemma 4.1 along with the resolvent estimate Eq. 4.4 with $\alpha=\beta=0$, we assert that
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\eta_{j}^{(0)}}\rVert_{L^{q}}^{p}\biggr{]}\text{ is uniformly bounded w.r.t.~{}$h$ and $\tau$ for all $p,q\in(2,\infty)$}.$ (4.31)
For any $p\in(2,\infty)$, the stability property of $P_{h}$ from Lemma 4.1 and the growth property of $F$ in Eq. 4.5 imply that
$\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\left(\lVert{Y_{j}-P_{h}Y_{j}^{3}}\rVert_{L^{2}}^{p}+\lVert{P_{h}F(Y_{j})}\rVert_{\gamma(H,L^{6})}^{p}\right)\biggr{]}$ $\leqslant c+c\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\lVert{Y_{j}}\rVert_{L^{6}}^{3p}\biggr{]}$
Moreover, it can be established that
$\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\lVert{Y_{j}}\rVert_{L^{6}}^{3p}\biggr{]}$ $\leqslant c\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\left(\lVert{Y_{j}-y_{h}(t_{j})}\rVert_{L^{6}}^{3p}+\lVert{y_{h}(t_{j})}\rVert_{L^{6}}^{3p}\right)\biggr{]}$
$\leqslant c+c\lVert{y_{h}}\rVert_{C([0,T];L^{3p}(\Omega;L^{6}))}^{3p},$
where the second inequality is justified by the error estimate Eq. 4.22, the condition $\tau\leqslant h^{2}$, and the application of the inverse estimate $\lVert{u_{h}}\rVert_{L^{6}}\leqslant ch^{-1}\lVert{u_{h}}\rVert_{L^{2}}$ for all $u_{h}\in X_{h}$. Consequently, by appealing to the stability assertion Eq. 4.17, we conclude that
$\mathbb{E}\biggl{[}\tau\sum_{j=0}^{J}\left(\lVert{Y_{j}-P_{h}Y_{j}^{3}}\rVert_{L^{2}}^{p}+\lVert{P_{h}F(Y_{j})}\rVert_{\gamma(H,L^{6})}^{p}\right)\biggr{]}$
remains uniformly bounded with respect to $h$ and $\tau$ for all $p\in(2,\infty)$. Using this result in conjunction with Lemmas 4.5 and 4.1, and taking advantage of the continuous embeddings of $\dot{H}_{h}^{2,2}$ and $\dot{H}_{h}^{1,6}$ into $L^{q}$ for any $q\in[2,\infty)$, we are able to assert that
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\left(\|\eta_{j}^{(1)}\|_{L^{q}}^{p}+\|\eta_{j}^{(2)}\|_{L^{q}}^{p}\right)\biggr{]}$
is also uniformly bounded with respect to $h$ and $\tau$ for all $p,q\in(2,\infty)$. Combining this finding with the assertion in (4.31) and noting that $Y_{j}=\eta_{j}^{(0)}+\eta_{j}^{(1)}+\eta_{j}^{(2)}$ for $1\leqslant j\leqslant J$, we confirm the stability property (4.23).
4.3.3 Proof of the error estimate Eq. 4.24
Firstly, according to Eq. 4.25, for each $0\leqslant j\leqslant J$, we decompose the error $Y_{j}-y_{h}(t_{j})$ into three components:
$Y_{j}-y_{h}(t_{j})=\chi_{j}^{(0)}+\chi_{j}^{(1)}+\chi_{j}^{(2)}.$ (4.32)
The sequences $(\chi_{j}^{(0)})_{j=0}^{J}$, $(\chi_{j}^{(1)})_{j=0}^{J}$, and $(\chi_{j}^{(2)})_{j=0}^{J}$ are recursively defined as follows:
$\begin{cases}\chi^{(0)}_{j+1}-\chi^{(0)}_{j}=\tau\Delta_{h}\chi^{(0)}_{j+1}+\int_{t_{j}}^{t_{j+1}}\Delta_{h}\big{(}y_{h}(t_{j+1})-y_{h}(t)\big{)}+Y_{j}-y_{h}(t)\,\mathrm{d}t,\quad 0\leqslant j<J,\\ \chi_{0}^{(0)}=0,\end{cases}$
[…]
and
$\begin{cases}\chi^{(2)}_{j+1}-\chi^{(2)}_{j}=\tau\Delta_{h}\chi^{(2)}_{j+1}+P_{h}\int_{t_{j}}^{t_{j+1}}F(Y_{j})-F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant j<J,\\ \chi_{0}^{(2)}=0.\end{cases}$
Secondly, by using Lemma 4.5, the continuous embedding of $\dot{H}_{h}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$, and the inequalities Eq. 4.18, Eq. 4.19, and Eq. 4.22, we derive the following bound for $(\chi_{j}^{(0)})_{j=1}^{J}$:
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(0)}}\rVert_{L^{q}}^{p}\biggr{]}$ $\leqslant\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{L^{q}}^{p}+\lVert{Y_{j}-y_{h}(t)}\rVert_{L^{2}}^{p}\,\mathrm{d}t\biggr{]}$
$\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.33)
Thirdly, we proceed to estimate the sequence $(\chi_{j}^{(1)})_{j=1}^{J}$. Let us fix $p,q\in(2,\infty)$. By the stability of $P_{h}$ as stated in Lemma 4.1 and Hölder’s inequality, a direct computation yields
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\right\|_{L^{3/2}}^{p}\,\mathrm{d}t\Biggr{]}$
$\leqslant c\left[\mathbb{E}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|y_{h}(t)-Y_{j+1}\right\|_{L^{2}}^{2p}\,\mathrm{d}t\right]^{1/2}\left[\mathbb{E}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|y_{h}(t)\right\|_{L^{12}}^{4p}+\left\|Y_{j+1}\right\|_{L^{12}}^{4p}\,\mathrm{d}t\right]^{1/2}.$
By the stability properties (4.17) and (4.23), it follows that
$\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}\right\|_{L^{3/2}}^{p}\,\mathrm{d}t\biggr{]}$ $\leqslant c\left[\mathbb{E}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|y_{h}(t)-Y_{j+1}\right\|_{L^{2}}^{2p}\right]^{1/2}.$
Hence, by applying Lemma 4.5 and the continuous embedding of $\dot{H}_{h}^{2,3/2}$ into $L^{q}$, we deduce that
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(1)}}\rVert_{L^{q}}^{p}\biggr{]}$ $\leqslant c\left[\mathbb{E}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|y_{h}(t)-Y_{j+1}\right\|_{L^{2}}^{2p}\right]^{1/2}.$
Thus, by combining (4.19) and (4.22), we infer the following bound for $(\chi_{j}^{(1)})_{j=1}^{J}$:
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(1)}}\rVert_{L^{q}}^{p}\biggr{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.34)
Fourthly, using Proposition 4.1 and the continuous embedding of $\dot{H}_{h}^{1,2}$ into $L^{6}$, we obtain that, for any $p\in(2,\infty)$,
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(2)}}\rVert_{L^{6}}^{p}\biggr{]}$ $\leqslant c\mathbb{E}\biggl{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|F(Y_{j})-F(y_{h}(t))\right\|_{\gamma(H,L^{2})}^{p}\,\mathrm{d}t\biggr{]}.$
By the Lipschitz continuity of $F$ as stated in (4.7), along with the inequalities (4.18) and (4.22), we establish the following bound for $(\chi_{j}^{(2)})_{j=1}^{J}$:
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(2)}}\rVert_{L^{6}}^{p}\biggr{]}$ $\leqslant c\tau^{p/2},\quad\forall p\in(2,\infty).$
Combining this bound with (4.32), (4.33), and (4.34) yields the error estimate
$\mathbb{E}\biggl{[}\tau\sum_{j=1}^{J}\left\|Y_{j}-y_{h}(t_{j})\right\|_{L^{6}}^{p}\biggr{]}\leqslant c\tau^{p/2},\quad\forall p\in(2,\infty).$
With this error estimate, the bound for $(\chi_{j}^{(2)})_{j=1}^{J}$ can be refined to
$\mathbb{E}\bigg{[}\tau\sum_{j=1}^{J}\lVert{\chi_{j}^{(2)}}\rVert_{L^{q}}^{p}\bigg{]}\leqslant c\tau^{p/2},\quad\forall p,q\in(2,\infty).$ (4.35)
Finally, considering the decomposition (4.32) and the identity $Y_{0}=y_{h}(0)$, we obtain the desired error estimate (4.24) by combining the bounds established in (4.33), (4.34), and (4.35).
4.3.4 Proof of the error estimate Eq. 4.16
Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. Let the sequences $(\chi_{j}^{(0)})_{j=0}^{J}$, $(\chi_{j}^{(1)})_{j=0}^{J}$, and $(\chi_{j}^{(2)})_{j=0}^{J}$ be constructed as in Subsection 4.3.3. By their construction, for any $1\leqslant j\leqslant J$, the $L^{q}$-norm of $\chi_{j}^{(0)}$ can be estimated almost surely as follows:
$\lVert{\chi_{j}^{(0)}}\rVert_{L^{q}}$ $\leqslant\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\Big{\lVert}{(I-\tau\Delta_{h})^{k-j}\big{[}\Delta_{h}\left(y_{h}(t_{k+1})-y_{h}(t)\right)+Y_{k}-y_{h}(t)\big{]}}\Big{\rVert}_{L^{q}}\,\mathrm{d}t$
$\leqslant\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{(I-\tau\Delta_{h})^{k-j}}\rVert_{\mathcal{L}(\dot{H}_{h}^{1/2+1/p,q},\dot{H}_{h}^{2,q})}\lVert{y_{h}(t)-y_{h}(t_{k+1})}\rVert_{\dot{H}_{h}^{1/2+1/p,q}}\,\mathrm{d}t$
$\quad{}+\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{(I-\tau\Delta_{h})^{k-j}}\rVert_{\mathcal{L}(\dot{H}_{h}^{0,q})}\lVert{Y_{k}-y_{h}(t)}\rVert_{L^{q}}\,\mathrm{d}t.$
Applying the resolvent estimate Eq. 4.4 and employing Hölder’s inequality, we obtain, for any $1\leqslant j\leqslant J$, almost surely,
$\lVert{\chi_{j}^{(0)}}\rVert_{L^{q}}$ $\leqslant c\left[\sum_{k=0}^{j-1}\int_{t_{k}}^{t_{k+1}}\lVert{y_{h}(t)-y_{h}(t_{k+1})}\rVert_{\dot{H}_{h}^{1/2+1/p,q}}^{p}+\lVert{Y_{k}-y_{h}(t)}\rVert_{L^{q}}^{p}\,\mathrm{d}t\right]^{\frac{1}{p}}.$
Consequently, for the sequence $(\chi_{j}^{(0)})_{j=1}^{J}$, we derive the following bound almost surely:
$\max_{1\leqslant j\leqslant J}\lVert{\chi_{j}^{(0)}}\rVert_{L^{q}}^{p}\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{y_{h}(t)-y_{h}(t_{j+1})}\rVert_{\dot{H}_{h}^{1/2+1/p,q}}^{p}+\lVert{Y_{j}-y_{h}(t)}\rVert_{L^{q}}^{p}\,\mathrm{d}t.$
Similarly, for $(\chi_{j}^{(1)})_{j=1}^{J}$, we have, almost surely,
$\max_{1\leqslant j\leqslant J}\lVert{\chi_{j}^{(1)}}\rVert_{L^{q}}^{p}\leqslant c\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\lVert{P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\rVert_{L^{q}}^{p}\,\mathrm{d}t.$
Combining these results leads to
$\mathbb{E}\left[\max_{1\leqslant j\leqslant J}\lVert{\chi_{j}^{(0)}+\chi_{j}^{(1)}}\rVert_{L^{q}}^{p}\right]$
[…]
by Eqs. 4.20, 4.18 and 4.24. Furthermore, utilizing the stability property of $P_{h}$, the error estimate Eq. 4.24, the inequality Eq. 4.19, and the stability results in Eq. 4.17 and Eq. 4.23, we can establish the following inequality
$\mathbb{E}\bigg{[}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\big{\lVert}{P_{h}y_{h}^{3}(t)-P_{h}Y_{j+1}^{3}}\big{\rVert}_{L^{q}}^{p}\,\mathrm{d}t\bigg{]}\leqslant c\tau^{p/2}.$
Therefore, we conclude that
$\Big{\lVert}{\max_{1\leqslant j\leqslant J}\big{\lVert}{\chi_{j}^{(0)}+\chi_{j}^{(1)}}\big{\rVert}_{L^{q}}}\Big{\rVert}_{L^{p}(\Omega)}\leqslant c\tau^{1/2}.$ (4.36)
Next, for the sequence $(\chi_{j}^{(2)})_{j=1}^{J}$, by the stability estimate established in Proposition 4.2, we deduce that
$\left\|\max_{1\leqslant j\leqslant J}\|\chi_{j}^{(2)}\|_{L^{q}}\right\|_{L^{p}(\Omega)}$ $\leqslant c\bigg{[}\mathbb{E}\sum_{j=0}^{J-1}\int_{t_{j}}^{t_{j+1}}\left\|P_{h}\big{(}F(y_{h}(t))-F(Y_{j})\big{)}\right\|_{\gamma(H,L^{q})}^{p}\,\mathrm{d}t\bigg{]}^{1/p}.$
Using the stability property of $P_{h}$ from Lemma 4.1 and the Lipschitz continuity of $F$ given by Eq. 4.7, and applying the results from Eq. 4.18 and Eq. 4.24, we further establish the following bound:
$\Big{\|}\max_{1\leqslant j\leqslant J}\|\chi_{j}^{(2)}\|_{L^{q}}\Big{\|}_{L^{p}(\Omega)}\leqslant c\tau^{1/2}.$ (4.37)
Finally, given the decomposition Eq. 4.32, the desired error estimate Eq. 4.16 follows by combining the bounds Eq. 4.36 and Eq. 4.37. This completes the proof of Theorem 3.1.
5 Numerical results
This section is dedicated to the numerical validation of the theoretical findings presented in this paper. The numerical experiments are conducted using the following model equation:
$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+y(t)\,\mathrm{d}\beta(t),\quad 0<t\leqslant T,\\ y(0)=v.\end{cases}$
Here, $\beta$ represents a given Brownian motion, the spatial domain $\mathcal{O}$ is $(0,1)^{3}$, and the initial condition is specified by the function $v$ defined as:
$v(x):=\sin(\pi x_{1})\sin(\pi x_{2})\sin(\pi x_{3})\quad\text{for all }x:=(x_{1},x_{2},x_{3})\in\mathcal{O}.$
To verify the spatial accuracy as established in Theorem 3.1, we set the final time to $T=0.01$ and the time step to $\tau=1\times 10^{-6}$. We then use the numerical solution computed with a spatial mesh size of $h=\frac{1}{2^{8}}$ as our reference solution. Furthermore, we perform simulations across $500$ independent paths. The numerical results presented in Figure 1 agrees well with the expected spatial convergence rate of $O(h^{2-\epsilon})$, as outlined in Eq. 3.2, for the cases of $p=q=2$, $p=q=4$, and $p=q=16$.
To further affirm the temporal accuracy claimed in Theorem 3.1, we set the terminal time to $T=0.1$ and adjust the spatial mesh size $h$ to meet the condition $h=\tau^{1/2}$, ensuring that the temporal discretization error is predominant. We use the numerical solution with a time step of $\tau=T/2^{13}$ as the reference solution. As before, we conduct simulations over $500$ independent paths. The numerical results displayed in Figure 2 confirm the temporal convergence rate of $O(\tau^{1/2})$, consistent with the theoretical convergence rate specified in Eq. 3.2.
In our theoretical framework, it is essential that the spatial domain $\mathcal{O}$ is a bounded, convex subset of $\mathbb{R}^{3}$ with a boundary of class $\mathcal{C}^{2}$. For the numerical experiments conducted in this study, we have chosen a cubic domain due to its simplicity and practicality, which facilitates both computational implementation and the interpretation of results. It is important to note that the theory of discrete stochastic maximal $L^{p}$-regularity remains an area of ongoing research.
Figure 1: $\tau=1.0\times 10^{-6}$, $T=0.01$,
Figure 2: $T=0.1$, $h=\tau^{1/2}$
6 Conclusions
In this paper, we have utilized discrete deterministic and stochastic maximal $L^{p}$ regularity estimates to derive a pathwise uniform convergence rate of $O\big{(}h^{2-\epsilon}+\tau^{1/2}\big{)}$ for a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. This convergence estimate employs general spatial $L^{q}$-norms and imposes mild conditions on the diffusion coefficient. Our work introduces a new approach to establishing pathwise uniform convergence within the context of general spatial $L^{q}$-norms for finite element-based full approximations of nonlinear stochastic parabolic equations.